### v1

# Lower Bounds for the Large Deviations of Selberg’s Central Limit Theorem

Abstract.
Let $\delta>0$ and $\sigma=\frac{1}{2}+\tfrac{\delta}{\log T}$. We prove that, for any $\alpha>0$ and $V\sim\alpha\log\log T$ as $T\to\infty$,
$\frac{1}{T}\text{meas}\big{\{}t\in[T,2T]:\log|\zeta(\sigma+{\rm i}\tau)|>V\big{\}}\geq C_{\alpha}(\delta)\int_{V}^{\infty}\frac{e^{-y^{2}/\log\log T}}{\sqrt{\pi\log\log T}}{\rm d}y,$
where $\delta$ is large enough depending on $\alpha$. The result is unconditional on the Riemann hypothesis. As a consequence, we recover the sharp lower bound for the moments on the critical line proved by Heap & Soundararajan and Radziwiłł & Soundararajan. The constant $C_{\alpha}(\delta)$ is explicit and is compared to the one conjectured by Keating & Snaith for the moments.
1. Main Results
The Riemann zeta function is defined on the half-plane ${\rm Re}\ s>1$ by
$\zeta(s)=\sum_{n\geq 1}n^{-s}=\prod_{p\ \text{prime}}(1-p^{-s})^{-1}.$
It can be continued to a meromorphic function on the whole complex plane $\mathbb{C}$ with a pole at $s=1$. The Riemann hypothesis states that all non-trivial zeros of the function lie on the critical line ${\rm Re}\ s=1/2$. This paper is concerned with the distribution of large values of the function on the critical line and slightly to its right.
A good point of reference for the distribution of values of the function on the line is Selberg’s Central Limit Theorem: if $\tau$ is a random point sampled uniformly in $[T,2T]$, then we have the following convergence in distribution
(1) $\frac{\log|\zeta(1/2+{\rm i}\tau)|}{\frac{1}{2}\sqrt{\log\log T}}\to\mathcal{N}(0,1),\quad T\to\infty,$
for $\mathcal{N}(0,1)$, a standard Gaussian random variable. It is not hard to see from the proof, see for example [RS17], that the result also holds slightly off-axis, i.e. ${\rm Re}\ s=\frac{1}{2}+\frac{\delta}{\log T}$, $\delta>0$. An important question, because of its relation to the moments of $\zeta$ and to the Lindelöf hypothesis, is to see if the distribution of the values of $\log|\zeta|$ remains Gaussian-like for values of the order of $\log\log T$, thus well beyond the standard deviation $\sqrt{\log\log T}$. It was established in [AB23] for $V\sim\alpha\log\log T$, $\alpha\in(0,2)$, that
(2) $\mathbf{P}(\log|\zeta(1/2+{\rm i}\tau)|>V)\leq K_{\alpha}\int_{V}^{\infty}\frac{e^{-y^{2}/\log\log T}}{\sqrt{\pi\log\log T}}{\rm d}y,$
for some constant $K_{\alpha}$ that depends on $\alpha$ but not $T$, and where $\mathbf{P}$ stands for the uniform probability on $\tau$. This in turn implied another proof for a sharp upper bound for the fractional moments of $\zeta$ between 00 and $4$ as first proved in [HRS19].
Here, we prove an unconditional lower bound off-axis matching (2) up to constant:
Theorem 1.
Let $\alpha>0$ and $V\sim\alpha\log\log T$. Consider $\tau$ sampled uniformly on $[T,2T]$. There exists $\delta^{\star}(\alpha)$ such that if $\delta>\delta^{\star}(\alpha)$, we have
$\mathbf{P}\Big{(}\log|\zeta(1/2+\delta/\log T+{\rm i}\tau)|>V\Big{)}\geq C_{\alpha}(\delta)\int_{V}^{\infty}\frac{e^{-y^{2}/\log\log T}}{\sqrt{\pi\log\log T}}{\rm d}y,$
for some $C_{\alpha}(\delta)>0$.
The constant $C_{\alpha}(\delta)$ is explicit, see Equation (19). So is $\delta^{\star}$, as given in Equation (33). For $\alpha$ tending to $\infty$, $\delta^{\star}(\alpha)$ is of order $\log\alpha$, whereas $\delta^{\star}(\alpha)$ remains bounded away from 00 as $\alpha\to 0$. (For technical reason, $\delta$ is redefined in (13), but this does not change the asymptotic behavior.) Our method of proof does not seem to apply at $\delta=0$, although it does imply sharp lower bound on the moments on the critical line. More precisely, it was conjectured by Keating & Snaith [KS00] that for all $k\geq 0$
(3) $\frac{1}{T}\int_{T}^{2T}|\zeta(1/2+{\rm i}t)|^{2k}{\rm d}t=\mathbf{E}\left[|\zeta(1/2+{\rm i}\tau)|^{2k}\right]\sim C_{k}\ (\log T)^{k^{2}},$
where $C_{k}=a_{k}f_{k}$. The arithmetic part $a_{k}$ is of the form
(4) $a_{k}=\prod_{p}(1-p^{-1})^{k^{2}}\sum_{m\geq 0}\left(\frac{\Gamma(k+m)}{m!\Gamma(k)}\right)^{2}p^{-m},$
(see also Equation (115)), and the random matrix part is given by
(5) $f_{k}=\frac{G^{2}(1+k)}{G(1+2k)},$
where $G$ is the Barnes $G$-function. Theorem 1 yields:
Corollary 1.
For $k\geq 0$ fixed, we have
$\mathbf{E}\left[|\zeta(1/2+{\rm i}\tau)|^{2k}\right]\geq c_{k}(\log T)^{k^{2}},$
where the constant $c_{k}$ is proportional to
(6) $a_{k}e^{\gamma k^{2}}(k+1)^{-38k^{2}}.$
Sharp lower bounds of this kind were proved in [HS22] and [RS13]. The constant obtained by Heap & Sound is better for large $k$ as it behaves like $k^{-2k^{2}}$. For $k$ approaching 00, their constant tends to 00 whereas the one in Corollary 1 is bounded away from it, as it should. Although, as mentioned in their paper, a modification of their argument would lead to the correct behavior for small $k$. Note that the constant $c_{k}$ includes explicitly the arithmetic part $a_{k}$, and this is in fact necessary in our argument. However, this does not change the asymptotic behavior for large $k$, since as shown in Conrey & Gonek [CG01], $\log a_{k}\sim-k^{2}\log\log k$ as $k\to\infty$. The correct asymptotic behavior of $f_{k}$ is $\log f_{k}\sim-k^{2}\log k$. It is unclear to us at this time how to improve the constant $c_{k}$, though a better approximation of indicator functions in terms of polynomials seems necessary, cf. Lemma 2 and the discussion following Equation (19). It is important to note that the same constant $C_{k}$ as for the moments should appear in the large deviation of $\log|\zeta|$. Indeed, it was conjectured by Radziwiłł that for any $\alpha>0$ and $V\sim\alpha\log\log T$ as $T\to\infty$,
(7) $\mathbf{P}(\log|\zeta(1/2+{\rm i}\tau)|>V)\sim C_{\alpha}\int_{V}^{\infty}\frac{e^{-y^{2}/\log\log T}}{\sqrt{\pi\log\log T}}{\rm d}y.$
The paper is organized as follows. The proof of Theorem 1 is explained in Section 2. It is divided in four propositions that are proved in Section 3. Corollary 1 is proved in Section 4. Finally, the appendix contains results needed in the proof that have appeared elsewhere in the literature. One notable exception is the mollification result stated in Lemma 5 that may be of independent interest. It was developed with one of the authors and Bourgade & Radziwiłł in an earlier version of [ABR23].
Notation. We write $f(T)=\operatorname{O}(g(T))$ whenever $\limsup_{T\to\infty}|f(T)/g(T)|<\infty$ with other parameters fixed, such as $\alpha$ and $k$. We also use Vinogradov’s notation $f(T)\ll g(T)$ whenever $f(T)=\operatorname{O}(g(T))$, and $f(T)\asymp g(T)$ if $f(T)\ll g(T)$ and $f(T)\gg g(T)$.
Acknowledgement. The authors are grateful to Winston Heap for his comments on the first version of this paper. L.-P. A. thanks Paul Bourgade and Maksym Radziwiłł for numerous insightful discussions on the subject, especially on the proof of Lemma 5. L.-P. A. is supported in part by the NSF award DMS 2153803.
2. Structure of the Proof of Theorem 1
2.1. Some definitions
As in [AB23], the proof relies on approximating $\log|\zeta|$ by Dirichlet polynomials of the form:
(8) $S_{\ell}=\sum_{T_{0}<p\leq T_{\ell}}\frac{{\rm Re}\ p^{-{\rm i}\tau}}{p^{\sigma}}+\frac{{\rm Re}\ p^{-2{\rm i}\tau}}{2p^{2\sigma}},\quad\ell>0.$
Throughout we use the probability convention of dropping $\tau$ in the notation. Also, $\sigma$ will be fixed and its dependence often omitted. The small primes need to be treated with care to get a good estimate on the constant. For the primes below $T_{0}$, we consider $|\mathcal{M}_{0}^{-1}|$ where
(9) $\mathcal{M}_{0}=\mathcal{M}_{0}(\sigma+{\rm i}\tau)=\prod_{p\leq T_{0}}\Big{(}1-p^{-\sigma-{\rm i}\tau}\Big{)}=\sum_{\begin{subarray}{c}p|m\Rightarrow p\leq T_{0}\end{subarray}}\frac{\mu(m)}{m^{\sigma+{\rm i}\tau}}.$
The parameter $T_{0}$ cannot be too large to ensure that $\mathcal{M}_{0}$ stays relatively short. An admissible choice turns out to be:
(10) $T_{0}=\log\log T.$
The parameters $T_{\ell}$ are chosen to approach $T$ fairly quickly:
[…]
The constant $\mathcal{L}$ is the largest $\ell$ for which $\log_{\ell+1}T>e$. By definition, we thus have $e<\log_{\mathcal{L}+1}T\leq e^{2}$ and $1<\log_{\mathcal{L}+2}T\leq e$. The parameter $\mathfrak{s}$ plays an important role in the analysis to estimate the constant $C_{\alpha}(\delta)$. The technical analysis requires $\mathfrak{s}=\mathfrak{s}(\alpha)$ to be not too small, to be precise:
(12) $\frac{(\log_{\mathcal{L}+1}T)^{\mathfrak{s}}}{(\mathfrak{s}\log_{\mathcal{L}+2}T)^{35}}>K(\alpha+1)^{35},$
where $K$ is a large absolute constant independent of $\alpha$. The power of $35$ is imposed during the proof of Proposition 2 (in particular in Lemma 2), see (55) and the surrounding discussion for further details.
It is convenient for the analysis to parametrize $\sigma>1/2$ in the following form:
(13) $\sigma=\frac{1}{2}+\frac{\delta}{\log T_{\mathcal{L}}}.$
2.2. Heuristic
By informally expanding the Euler product in the definition of $\zeta$, we expect that $\log|\zeta|$ at $\sigma+{\rm i}\tau$ should behave like
$\log|\zeta(\sigma+{\rm i}\tau)|\approx S_{\mathcal{L}}+\log|\mathcal{M}_{0}|^{-1}.$
Moreover, since $S_{\ell}$ is a sum of weakly correlated terms, it should behave like a Gaussian random variable for large $T$. With this in mind, we define the random variables
(14) $\mathcal{Z}_{\ell}=\sum_{T_{0}<p\leq T_{\ell}}\frac{Z_{p}}{p^{\sigma}}+\frac{Z_{p}^{2}}{2p^{2\sigma}},\quad 1\leq\ell\leq\mathcal{L}.$
We effectively replaced $p^{-{\rm i}\tau}$ by $Z_{p}$ where $(Z_{p},\text{p prime})$ are centered Gaussian random variables with variance $1/2$. The contribution of very small primes is not Gaussian even in the limit $T\to\infty$. Instead, for the small primes, we substitute the phase $p^{-{\rm i}\tau}$ in by $e^{{\rm i}\theta_{p}}$, where $(\theta_{p},\text{p prime})$ are IID uniform on $[0,2\pi]$:
(15) $\mathcal{Z}_{0}=\log\prod_{p\leq T_{0}}\Big{|}1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\Big{|}^{-1}=-\sum_{p\leq T_{0}}\sum_{k\geq 1}\frac{\cos k\theta_{p}}{kp^{k\sigma}}.$
We thus expect for $V\sim\alpha\log\log T$, $\alpha>0$, that
(16) $\mathbf{P}(\log|\zeta(\sigma+{\rm i}\tau)|>V)\approx\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
The variance of $\mathcal{Z}_{\mathcal{L}}$ is not exactly $1/2(\log\log T_{\ell}-\log\log T_{0})$ as one would expect from Mertens’s theorem, since $\sigma>1/2$. In fact, we have for any $1\leq\ell\leq\mathcal{L}$ by the Prime Number Theorem
$v_{\ell}^{2}=\mathbf{E}[\mathcal{Z}_{\ell}^{2}]=\sum_{T_{0}<p\leq T_{\ell}}\Big{(}\frac{1}{2p^{2\sigma}}+\frac{1}{8p^{4\sigma}}\Big{)}$ $=\frac{1}{2}\int_{\log T_{0}}^{\log T_{\ell}}\frac{e^{-(2\sigma-1)u}}{u}{\rm d}u+\operatorname{O}(e^{-c\sqrt{\log T_{0}}})$
(17) $=\frac{1}{2}\left(\log\log T_{\ell}-\log\log T_{0}\right)-\delta\frac{\log T_{\ell}}{\log T_{\mathcal{L}}}+\operatorname{O}(e^{-c\sqrt{\log T_{0}}}).$
The conclusion of Theorem 1 is then simply the result of a Gaussian estimate, conditioning on $\mathcal{Z}_{0}$, and of expanding $v_{\mathcal{L}}$:
$\mathbf{P}(\mathcal{Z}_{\mathcal{L}}+\mathcal{Z}_{0}>V)$ $\asymp\mathbf{E}\Big{[}\frac{e^{-(V-\mathcal{Z}_{0})^{2}/(2v_{\mathcal{L}}^{2})}}{\alpha\sqrt{\log\log T}}\Big{]}$
(18) $\asymp\left(\frac{\mathbf{E}[e^{2\alpha\mathcal{Z}_{0}}]}{(\log T_{0})^{\alpha^{2}}}\cdot e^{-2\alpha^{2}\delta}\cdot(\log_{\mathcal{L}+1}T)^{-\alpha^{2}\mathfrak{s}}\right)\frac{e^{-V^{2}/\log\log T}}{\alpha\sqrt{\log\log T}}\ .$
The term $(\log T_{0})^{-\alpha^{2}}\mathbf{E}[e^{2\alpha\mathcal{Z}_{0}}]$ turns out to be exactly the arithmetic factor $e^{\gamma\alpha^{2}}a_{\alpha}$ by a short computation on the random model, cf. Lemma 12. We therefore obtain the lower bound claimed in Theorem 1 with
(19) $C_{\alpha}(\delta)=e^{\gamma\alpha^{2}}a_{\alpha}\cdot e^{-2\alpha^{2}\delta}\cdot(\alpha+1)^{-36\alpha^{2}},$
by taking $(\log_{\mathcal{L}+1}T)^{\mathfrak{s}}=K(\alpha+1)^{36}$, which satisfies the bound (12). The significance of each factor constituting $C_{\alpha}(\delta)$ is clear from the heuristic. The arithmetic factor $a_{\alpha}$ is produced by the small primes. The term $e^{-2\alpha^{2}\delta}$ is due to the shift off-axis and should be 00 on the critical line (perhaps by assuming the Riemann hypothesis). Finally, the factor $(\alpha+1)^{-36\alpha^{2}}$ is due to the cutoff of the large prime at $T^{1/(\log_{\mathcal{L}+1}T)^{\mathfrak{s}}}$. The optimal cutoff should be expected to be $T^{c/\alpha}$ for some small constant $c>0$. Such a cutoff would yield a factor of $(\alpha+1)^{-\alpha^{2}}$ corresponding to the asymptotics of $f_{\alpha}$ in (5) predicted by random matrix theory.
2.3. Proof of Theorem 1
We now outline the roadmap to establish a rigorous lower bound for (16). An approximation of $\log|\zeta|$ in terms of the partial sums $S_{\ell}$ and of the Gaussian sums $\mathcal{Z}_{\ell}$ can be made precise on a good event where the sums are constrained. Specifically, we define
(20) $G_{0}=\{|\mathcal{M}_{0}|^{2}\in[L_{0},U_{0}]\}.$
and
(21) $G_{\ell}=G_{\ell-1}\cap\{L_{\ell}(S_{0})\leq S_{\ell}\leq U_{\ell}(S_{0})\},1\leq\ell\leq\mathcal{L},$
where the upper and lower barriers $L_{\ell}(z),U_{\ell}(z)$ are defined and discussed below in (25), and $S_{0}=-\frac{1}{2}\log|\mathcal{M}_{0}|$, corresponding to the definition (9) and the choice in the initial good event $G_{0}$. (This event is chosen since $\mathcal{M}_{0}$ is a short Dirichlet polynomial, cf. (43).) We take
(22) $L_{0}=\exp(-2\alpha\log\log T_{0})\qquad U_{0}=\exp(-2\alpha\log\log T_{0}+2\sqrt{\log\log T_{0}}).$
The factor of $2$ in (22) is a direct consequence of taking the squared polynomial in (20). The final event of interest is
(23) $G(V)=\{\log|\mathcal{M}_{0}|^{-1}+S_{\mathcal{L}}>V\}\cap G_{\mathcal{L}}.$
We define the corresponding events for the sums $\mathcal{Z}_{\ell}$,
$\mathcal{G}_{0}$ =$=\{\mathcal{Z}_{0}-\alpha\log\log T_{0}\in[0,\sqrt{\log\log T_{0}}]\}$
[…]
and
(24) $\mathcal{G}(V)=\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\}\cap\mathcal{G}_{\mathcal{L}}.$
In $G_{0}$ and $\mathcal{G}_{0}$, the upper bound is natural, coming from the standard deviation in both cases. It is convenient to assume that the recentered sum in case is positive, and we use this fact in the proof of Proposition 1, cf. (35). The (random) upper and lower barriers $U_{\ell}(z)$ and $L_{\ell}(z)$ are defined by
(25) $L_{\ell}(z)$ $=m(z)(\log\log T_{\ell}-\log\log T_{0})-\mathcal{C}\log_{\ell+2}T$
$U_{\ell}(z)$ $=m(z)(\log\log T_{\ell}-\log\log T_{0})+\mathcal{B}\log_{\ell+2}T$
where the slope $m(z)$ is defined as
(26) $m(z)=\frac{\alpha\log\log T-z}{\log\log T_{\mathcal{L}}-\log\log T_{0}}.$
where $\log_{\ell}$ stands for the logarithm iterated $\ell$ times. In each case, the slope $m(z)$ will be evaluated at the (random) initial position, i.e. $S_{0}$ or $\mathcal{Z}_{0}$. Taking the barriers randomly is necessitated by separating out the contribution from the small primes. On first reading, one can think of the slope $m(z)\approx\alpha$ (since typically $z$ is like $\alpha\log\log T_{0}$ in each case). However, we need to take in to account the variance coming from the initial random sums which informs the slope depending on the random position of $\mathcal{Z}_{0}$ or $S_{0}$, see in particular the remark following (36). See also Figure 1, where two different initial values $\mathcal{Z}_{0}$ and $\mathcal{Z}^{\prime}_{0}$ both reach level $V$ but elicits two different slopes.
The constants $\mathcal{B}$ will need to be larger than $\alpha\mathfrak{s}$ and $\mathcal{C}$ a sufficiently large constant
(27) $\mathcal{B}=\alpha\mathfrak{s}+\frac{1}{\alpha}\qquad\mathcal{C}=100.$
We include the term $1/\alpha$ in $\mathcal{B}$ to allow for small $\alpha$, cf. Section 3.1.
[…]
Figure 1. Illustration of the random sums starting at two distinct initial values, both ending at the same point $V$. The lines $f_{z}(T_{j})$ have (distinct) slopes $m(z)$ for $z=\mathcal{Z}_{0},\mathcal{Z}^{\prime}_{0}$.
The rationale behind the choice of restrictions is that either sum $S_{\ell}$ or $\mathcal{Z}_{\ell}$, whose total value is $V\sim\alpha\log\log T$, should stay close to the linear interpolation with slope $m(z)$ with some fluctuations that are getting smaller with $\ell$. It is not hard to first establish this fact for the random model:
Proposition 1.
For fixed $\alpha>0$ and $V\sim\alpha\log\log T$, the probability of the event $\mathcal{G}(V)$ defined in (23) is
$\mathbf{P}(\mathcal{G}(V))\asymp\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
To prove this for the partial sums $S_{\ell}$ is much trickier. To do so, we follow the idea of [ABR20] and [ABR23] and approximate the events involving $S_{\ell}$ and $\mathcal{M}_{0}$ in terms of short Dirichlet polynomials. This allows us to explicitly compare the probability of the event $G(V)$ with the event $\mathcal{G}(V)$ of the random model. We ultimately show
Proposition 2.
For fixed $\alpha>0$ and $V\sim\alpha\log\log T$, the probability of the event $\mathcal{G}(V)$ defined in (24) is
(28) $\mathbf{P}(G(V))\asymp\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
It remains to connect the event $G(V)$ to $\log|\zeta|$. This is done by defining the following short mollifiers:
(29) $\mathcal{M}_{\ell}=\sum_{\begin{subarray}{c}p|m\Rightarrow p\in(T_{\ell-1},T_{\ell}]\\ \Omega_{\ell}(m)\leq\mu_{\ell}\end{subarray}}\frac{\mu(m)}{m^{\sigma+{\rm i}\tau}},\quad 1\leq\ell\leq\mathcal{L},$
where $\Omega_{\ell}(m)$ is the number of prime factors of $m$ in the interval $(T_{\ell-1},T_{\ell}]$, and we choose
(30) $\qquad\mu_{\ell}=100(\alpha\vee 1)(\log\log T_{\ell}-\log\log T_{\ell-1}),\quad 1\leq\ell\leq\mathcal{L}.$
This is to ensure the length of the mollifiers is not too long. Indeed, taking the product over $\ell$, we set (recalling the definition of $\mathcal{M}_{0}$ from (9))
(31) $\mathcal{M}=\prod_{0\leq\ell\leq\mathcal{L}}\mathcal{M}_{\ell}.$
Then, the definition of $\mu_{\ell}$ and the choice of $\mathfrak{s}$ in (12) implies that $\mathcal{M}$ is supported on integers smaller than
(32) $\prod_{0\leq\ell\leq\mathcal{L}}(T_{\ell})^{\mu_{\ell}}\ll T^{200(\alpha+1)\mathfrak{s}(\log_{\mathcal{L}+1}T)^{1-\mathfrak{s}}}\leq T^{\frac{1}{K(\alpha+1)^{34}}}<T^{1/100},$
for a sufficiently large $K$. The next proposition shows that $\zeta\mathcal{M}$ is close to $1$ on the event $G(V)$.
Proposition 3.
Let $\varepsilon\in(0,1)$. For fixed $\alpha>0$ and $V\sim\alpha\log\log T$, we have
$\mathbf{P}(\log|\zeta|>V)\gg\mathbf{P}\Big{(}\{\log|\mathcal{M}^{-1}|>V-\log(1-\varepsilon)\}\;\cap\;G(V)\Big{)}-\frac{1}{\varepsilon^{2}}\left(\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\right)\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
The proposition establishes an interplay between $\varepsilon$ and $\delta$: if $\varepsilon$ is close to 00 or $1$, then the expression could be negative. In the proof of Theorem 1, we find an optimal $\varepsilon$ and derive the threshold $\delta^{\star}$ from it. The last proposition relates the inverse of the mollifier to the partial sums, which can then be compared to $\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V)$ using Propositions 2 and 1.
Proposition 4.
For fixed $\alpha>0$ and $V\sim\alpha\log\log T$, we have for the event $G(V)$ in (23)
$\mathbf{P}\left(\{\log|\mathcal{M}^{-1}|>V\}\cap G(V)\right)\gg\mathbf{P}\left(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\right).$
The above propositions combine to give a proof of the theorem.
Proof of Theorem 1.
Consider $V$ with $V\sim\alpha\log\log T$. For some $\varepsilon\in(0,1)$ to be picked below, define $V^{\prime}=V-\log(1-\varepsilon)$. Then, Propositions 3 and 4 directly imply
$\mathbf{P}(\log|\zeta|>V)\gg\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V-\log(1-\varepsilon))-\frac{1}{\varepsilon^{2}}\left(\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\right)\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
We have from (19) that this is
$\gg C_{\alpha}(\delta)\frac{e^{-V^{2}/\log\log T}}{\alpha\sqrt{\log\log T}}\Big{\{}(1-\varepsilon)^{2\alpha}-\frac{1}{\varepsilon^{2}}\Big{(}\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\Big{)}\Big{\}}.$
For a fixed $\alpha$, we are looking for the smallest $\delta$ that makes the quantity in the parenthesis positive. The function $\varepsilon^{2}(1-\varepsilon)^{2\alpha}$ is maximal at $\varepsilon^{\star}=1/(\alpha+1)$. We thus pick $\delta^{\star}$ to be equal to a fraction (say $e^{-1}$) of $\varepsilon^{2}(1-\varepsilon)^{2\alpha}$:
$\frac{e^{-2\delta^{\star}}}{\delta^{\star}}=\frac{1}{e}{\varepsilon^{\star}}^{2}(1-\varepsilon^{\star})^{2\alpha}=\frac{1}{e}\frac{\alpha^{2\alpha}}{(\alpha+1)^{2(\alpha+1)}},$
which gives
(33) $\delta^{\star}+\frac{1}{2}\log\delta^{\star}=\log(\alpha+1)+\alpha\log(\alpha^{-1}+1)+\frac{1}{2}.$
Since $\frac{e^{-2\delta}}{\delta}$ is a decreasing function of $\delta$ and that $(1-\varepsilon^{\star})^{2\alpha}\asymp 1$, we get that for $\delta>\delta^{\star}$ the above is
$\gg C_{\alpha}(\delta)\frac{e^{-V^{2}/\log\log T}}{\alpha\sqrt{\log\log T}}\Big{(}\frac{\alpha}{\alpha+1}\Big{)}^{2\alpha}\gg C_{\alpha}(\delta)\frac{e^{-V^{2}/\log\log T}}{\alpha\sqrt{\log\log T}}.$
The term $e^{-\mu_{\mathcal{L}}}$ does not contribute by the choice of $\mu_{\ell}$ in (30) by taking $\mathfrak{s}$ large enough independently of $\alpha$, as mentioned below (12). ∎
3. Proof of the Propositions
Throughout the proofs, we use the following notations for conciseness:
$t=\log\log T,$
and
[…]
3.1. Proof of Proposition 1
The upper bound is immediate by dropping the event $\mathcal{G}_{\mathcal{L}}$, so we focus on the lower bound. Decomposing the event $\mathcal{G}_{\mathcal{L}}$ we inductively have
(34) $\mathbf{P}(\mathcal{G}(V))$ $=\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\}\cap\mathcal{G}_{0})-\sum_{\ell=0}^{\mathcal{L}-1}\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\}\cap\mathcal{G}_{\ell}\cap\mathcal{G}_{\ell+1}^{c}).$
We subsequently bound the summands on the right-hand side of (34) for $V\sim\alpha\log\log T=\alpha t$. Separately considering the two possibilities for $\mathcal{G}_{\ell+1}^{c}$, we have for the upper barrier
(35) $\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\}\cap\mathcal{G}_{\ell}\cap\{\mathcal{Z}_{\ell+1}>U_{\ell+1}(\mathcal{Z}_{0})\})$ $\leq\mathbf{P}\{\mathcal{Z}_{\ell+1}+\mathcal{Z}_{0}>U_{\ell+1}(\alpha t_{0})+\alpha t_{0}\})$
$\asymp\frac{1}{\alpha\sqrt{t}}\mathbf{E}\left[\exp\left(-\frac{(U_{\ell+1}(\alpha t_{0})+\alpha t_{0}-\mathcal{Z}_{0})^{2}}{2v_{\ell+1}^{2}}\right)\right]$
akin to (18). The first bound uses positivity of the recentered sum. Unfolding and using the definition of $U_{\ell+1}(z)$ in (25), we have that the above is bounded by
$\ll\frac{1}{\alpha\sqrt{t}}\mathbf{E}\left[\exp\left(-\frac{1}{2v_{\ell+1}^{2}}\left(\frac{\alpha}{t_{\mathcal{L}}-t_{0}}(t_{\ell+1}\left(t-t_{0})-t_{0}(t-t_{\mathcal{L}})\right)+\mathcal{B}\log_{\ell+1}t-\mathcal{Z}_{0}\right)^{2}\right)\right]$
$\ll\frac{1}{\alpha\sqrt{t}}e^{-\alpha^{2}t}e^{-\alpha^{2}t_{0}}\mathbb{E}\left[e^{2\alpha\mathcal{Z}_{0}}\right]e^{-2\alpha^{2}\delta e^{t_{\ell+1}-t_{\mathcal{L}}}-2\alpha^{2}\mathfrak{s}\log_{\mathcal{L}}t}\cdot e^{(\alpha^{2}\mathfrak{s}-2\alpha\mathcal{B})\log_{\ell+1}t}$
$\ll\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t)\cdot e^{(\alpha^{2}\mathfrak{s}-2\alpha\mathcal{B})\log_{\ell+1}t},$
the final bound achieved by using Lemma 12 and comparing to (18). The coefficient in the exponent involving $\log_{\ell+1}t$ is strictly negative by choosing $\mathcal{B}=\alpha\mathfrak{s}+1/\alpha$ (cf. Equation (27)), and the implicit constant can be taken independent of $\alpha$.
To complete the bound, we estimate the exceedance of the lower barrier
(36) $\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t\}\cap\mathcal{G}_{0}\cap\{\mathcal{Z}_{\ell+1}<L_{\ell+1}(\mathcal{Z}_{0})\}).$
If one uses a deterministic barrier $L_{\ell+1}\approx\alpha(t_{\ell+1}-t_{0})-\mathcal{C}\log_{\ell+1}t$ here (corresponding to the typical position of $\mathcal{Z}_{0}$) then the position of the process at $\ell+1$ is at most $L_{\ell+1}+\mathcal{Z}_{0}$. This position could be as large as $\approx\alpha t_{\ell+1}+\sqrt{t_{0}}+\mathcal{C}\log_{\ell+1}t$. This additional term coming from the variance of the initial contribution does not permit a strong enough bound in the calculation below.
First conditioning on the value of $\mathcal{Z}_{0}$ and decomposing on the value at $t_{\ell+1}$ we find:
$\mathbf{P}(\{\mathcal{Z}_{\mathcal{L}}-\mathcal{Z}_{\ell+1}$ $>\alpha t-\mathcal{Z}_{\ell+1}-w\}\cap\{\mathcal{Z}_{\ell+1}<L_{\ell+1}(w)\})$
$\leq\sum_{u\leq L_{\ell+1}(w)}\mathbf{P}(\mathcal{Z}_{\mathcal{L}}-\mathcal{Z}_{\ell+1}>\alpha t-(u+1)-w)\mathbf{P}(\mathcal{Z}_{\ell+1}\in(u,u+1]),$
using independence. By the Gaussian decay, this is therefore bounded by
$\ll\sum_{u\leq L_{\ell+1}(w)}e^{-\frac{(\alpha t-(u+1)-w)^{2}}{2(v_{\mathcal{L}}^{2}-v_{\ell+1}^{2})}}\frac{v_{\ell+1}}{u}e^{-\frac{u^{2}}{2v_{\ell+1}^{2}}}$
$\ll\frac{1}{\alpha\sqrt{t}}e^{-\alpha^{2}\frac{(t_{\mathcal{L}}-t_{\ell+1})^{2}}{2(v_{\mathcal{L}}^{2}-v_{\ell+1}^{2})}}e^{-\alpha^{2}\frac{t^{2}}{2v_{\ell+1}^{2}}}e^{2\alpha t\frac{\alpha(t_{\mathcal{L}}-t_{\ell+1})+w}{2v_{\ell+1}^{2}}}\sum_{u>\mathcal{C}\log_{\ell+1}t}e^{-2\alpha\frac{(u-1)(t_{\mathcal{L}}-t_{\ell+1})}{2(v_{\mathcal{L}}^{2}-v_{\ell+1}^{2})}}e^{-\frac{(u-1)^{2}}{2(v_{\mathcal{L}}^{2}-v_{\ell+1}^{2})}}e^{2\alpha t\frac{u}{2v_{\ell+1}^{2}}},$
achieved by shifting the sum by $(t_{\ell+1}-t_{0})m(w)$ and expanding. The pre-factor simplifies to
$\frac{1}{\alpha\sqrt{t}}\exp\left(-\alpha^{2}t\right)\cdot\exp\left((-\alpha^{2}t_{0}+2\alpha w)-2\alpha^{2}\delta-\alpha^{2}\mathfrak{s}\log_{\mathcal{L}}t\right).$
After unconditioning on the value of $\mathcal{Z}_{0}$, this term is like $\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t)$. Similarly, after simplifying and estimating at the dominant $u$, the sum is bounded by
$\ll\exp\left(2\alpha-4\alpha\mathcal{C}\delta\frac{\log_{\ell+1}t}{t_{\mathcal{L}}-t_{\ell+1}}(1-e^{t_{\ell+1}-t_{\mathcal{L}}})-\frac{(\mathcal{C}\log_{\ell+1}t-1)^{2}}{2(v_{\mathcal{L}}^{2}-v_{\ell+1}^{2})}\right)\ll c_{\alpha}\exp\left(-\frac{\mathcal{C}^{2}}{\mathfrak{s}}\log_{\ell+1}t\right).$
Again, the coefficient in the final exponent involving $\log_{\ell+1}t$ is strictly negative (and we implicitly use that $\mathcal{C}$ is strictly positive when estimating the sum above), resulting in a subdominant contribution for most $\ell$. For the largest [^1] For $\ell=\mathcal{L}-1$, the probability involved is zero since the events $\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t\}$ and $\mathcal{Z}_{\mathcal{L}}<L_{\mathcal{L}}(\mathcal{Z}_{0})$ are incompatible. $\ell$, namely $\mathcal{L}-2$, notice that $e\leq\log_{\ell+1}t<e^{2}$. The term $c_{\alpha}$ is an explicit coefficient dependent on $\alpha$,
$c_{\alpha}=\exp\left(2\alpha-\frac{\mathcal{C}}{\mathfrak{s}}\left(4\alpha\delta+\mathcal{C}\log_{\mathcal{L}}t+2\mathcal{C}\delta-2\right)\right).$
The coefficient will appear in the sum over $\ell$ below. Recall $\mathfrak{s}=35\log(\alpha+1)+\log(2\cdot 10^{6})$ and $\delta>\delta^{\star}(\alpha)$ (cf. (55) and (33) respectively). Therefore for small $\alpha$, both $\mathcal{C}/\mathfrak{s},2\delta>1$ so $c_{\alpha}<1/100$ say. For large $\alpha$, we have $\delta^{\star}/\mathfrak{s}\approx 1$ so again $c_{\alpha}<\exp(-10\alpha)$ say.
Therefore, returning to the original probability, unconditioning and comparing to (18), we have for $20\leq\ell\leq\mathcal{L}-$ 2
$\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}$ $>\alpha t\}\cap\mathcal{G}_{0}\cap\{\mathcal{Z}_{\ell+1}<L_{\ell+1}(\mathcal{Z}_{0})\})\ll\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t)\cdot c_{\alpha}e^{-\frac{\mathcal{C}^{2}}{\mathfrak{s}}\log_{\ell+1}t}.$
Overall from (34) (using the estimate (117) to remove $\mathcal{G}_{0}$ in the first instance) we find
$\mathbf{P}(\mathcal{G}(V))$ $\gg\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V)\left(1-\sum_{j=0}^{\mathcal{L}-1}e^{-(2\alpha\mathcal{B}-\alpha^{2}\mathfrak{s})\log_{\ell+1}t}-c_{\alpha}\sum_{j=0}^{\mathcal{L}-2}e^{-\frac{\mathcal{C}^{2}}{\mathfrak{s}}\log_{\ell+1}t}\right)$
$\asymp\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V),$
where the final bound follows from the rapid convergence of the sum and taking $\mathcal{B},\mathcal{C}$ as in (27) (note that $c_{\alpha}$ is also small according to the discussion above) so that the term in brackets is strictly positive and can be taken independent of $\alpha$.
3.2. Proof of Proposition 2
We first need two lemmas whose goal is to approximate quantitatively an indicator function by a polynomial. The content of these lemmas can be found in Lemma 3 in [ABR23]. We reproved them here for completeness, and also to get a clear dependence on the parameters, which is key in estimating the constant $C_{k}$.
Lemma 1.
Let $c>b>a>1$ and $\Delta$ large enough. There exist entire functions $h^{+}$ and $h^{-}$ depending on these parameters such that for all $x\in\mathbb{R}$,
[…]
    $0\leq h^{-}\leq h^{+}\leq 1,$
[…]
    $\mathbf{1}(x\in[0,\Delta^{-1}])\Big{(}1-e^{-\Delta^{c-b-1}}\Big{)}\leq h^{+}(x)\leq\mathbf{1}(x\in[-\Delta^{-a},\Delta^{-1}+\Delta^{-a}])+e^{-\Delta^{c-a-1}},$
[…]
    $\mathbf{1}(x\in[\Delta^{-a},\Delta^{-1}-\Delta^{-a}])\Big{(}1-e^{-\Delta^{c-a-1}}\Big{)}\leq h^{-}(x)\leq\mathbf{1}(x\in[0,\Delta^{-1}])+e^{-\Delta^{c-b-1}},$
[…]
    for all $\ell\in\mathbb{N}$, the Fourier transform $\widehat{h^{\pm}}$ is such that
    $\left|\int_{\mathbb{R}}\xi^{\ell}\widehat{h^{\pm}}(\xi){\rm d}\xi\right|\leq\Delta^{c(\ell+2)}.$
Proof.
We define $h^{\pm}$ as convolutions with a suitable kernel. Following Lemma 5 in [ABR20] based on a classical construction of [Ing34], we know there exists a real function $F$ such that $0\leq F\leq 1$ with decay
$F(x)\ll\exp(-|x|/\log^{2}(|x|+10)),$
for which the Fourier transform is supported on $[-1,1]$. With this in mind, consider the corresponding PDF $\varphi(x)=F(x)/\|F\|_{1}$, and the scaled version:
$\varphi_{c}(y)=\Delta^{c}\varphi(\Delta^{c}y).$
Since $c>1$, the scale of $\varphi_{c}$ is much smaller than the width of the interval $[0,\Delta^{-1}]$ to be approximated. The function $h^{\pm}$ are defined by the convolution of an indicator function with $\varphi_{c}$:
(37) $h^{+}(x)=\int\mathbf{1}(y\in[-\Delta^{-b},\Delta^{-1}+\Delta^{-b}])\varphi_{c}(x-y){\rm d}y.$
and
(38) $h^{-}(x)=\int\mathbf{1}(y\in[\Delta^{-b},\Delta^{-1}-\Delta^{-b}])\varphi_{c}(x-y){\rm d}y.$
The functions $h^{\pm}$ should be thought of as smoothings of the indicator function $1(x\in[0,\Delta^{-1}])$. The slight overspill $\Delta^{-b}$ must be added to properly bound the indicator function close to 00 and $\Delta^{-1}$. It is necessary to take $c>b>1$, as this ensures that the effective width of the kernel is much smaller than the overspill. Moreover, to bound $h^{+}$ above (and $h^{-}$ below) by an indicator function, the interval has to be enlarged (or reduced) by $\Delta^{-a}$. For this, we need $b>a$ to ensure the indicator function in the definition of $h^{+}$ is bounded by $\mathbf{1}(x\in[-\Delta^{-a},\Delta^{-1}+\Delta^{-a}])$, and similarly for $h^{-}$.
Property i is obvious from the definitions (37) and (38). The first inequality in Property ii is also clear when $x\notin[0,\Delta^{-1}]$. The case $x\in[0,\Delta^{-1}]$ becomes
$1-e^{-\Delta^{c-b-1}}\leq h^{+}(x).$
To prove this, note that for any $0\leq x\leq\Delta^{-1}$,
$h^{+}(x)=\int_{-\Delta^{-1}-\Delta^{-b}}^{\Delta^{-b}}\varphi_{c}(u+x){\rm d}u\geq\int_{-\Delta^{-b}}^{\Delta^{-b}}\varphi_{c}(u){\rm d}u\geq 1-e^{-\Delta^{c-b-1}},$
where the last inequality follows by the decay of $\varphi$. The right-hand side inequality of Property ii is also clear when $x\in[-\Delta^{-a},\Delta^{-1}+\Delta^{-a}])$. If $x<-\Delta^{-a}$, then we have
$h^{+}(x)=\int_{-\Delta^{-1}-\Delta^{-b}}^{\Delta^{-b}}\varphi_{c}(u+x){\rm d}u\leq\int_{-\infty}^{\Delta^{-b}-\Delta^{-a}}\varphi_{c}(u){\rm d}u\leq e^{-\Delta^{c-a-1}},$
for $\Delta$ large enough and by the decay of $\varphi$ (since $b>a$, we can ask for $\Delta^{-b}-\Delta^{-a}\leq-\frac{1}{2}\Delta^{-a}$ say). Similarly, if $x>\Delta^{-1}+\Delta^{-a}$, we get
$h^{+}(x)=\int_{-\Delta^{-1}-\Delta^{-b}}^{\Delta^{-b}}\varphi_{c}(u+x){\rm d}u\leq\int_{\Delta^{-a}-\Delta^{-b}}^{\infty}\varphi_{c}(u){\rm d}u\leq e^{-\Delta^{c-a-1}}.$
Property iii is proved exactly as Property ii by using the definition of $h^{-}$ in (38) instead.
For Property iv, note that
$\widehat{h^{\pm}}(\xi)=\widehat{\varphi}_{c}(\xi)\cdot\widehat{\mathbf{1}_{[0,\Delta^{-1}]}}(\xi).$
On the one hand, by integration, we have the simple bound
$\widehat{\mathbf{1}_{[0,\Delta^{-1}}]}(\xi)\ll\Delta^{-1}.$
This smaller than $1$ for $\Delta$ large enough. On the other hand, since $\varphi$ is supported on $[-1,1]$, we have
$\left|\int\xi^{\ell}\widehat{\varphi}_{c}(\xi){\rm d}\xi\right|\leq\Delta^{c\ell}\int|\widehat{\varphi_{c}}(\xi)|{\rm d}\xi.$
By the Cauchy-Schwarz inequality and Parseval’s identity, we can bound the integral by
$\int|\widehat{\varphi_{c}}(\xi)|{\rm d}\xi\leq\sqrt{2\Delta^{c}}\left(\int\varphi_{c}(x)^{2}{\rm d}x\right)^{1/2}\ll\Delta^{c},$
since $\varphi_{c}(x)\ll\Delta^{c}$. This can be made $\leq\Delta^{2c}$ by taking $\Delta$ large enough to absorb the implicit constant. This gives the desired bound. ∎
In a second step, we approximate $h^{\pm}$ by polynomials.
Lemma 2.
Let $a>2$, $X>1$ and $\Delta$ large enough. There exist polynomials $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$ of degree smaller than $100X\Delta^{3a}$ with $\ell$-th coefficient bounded by $\frac{(2\pi)^{\ell}}{\ell!}\Delta^{2a(\ell+2)}$ such that for $|x|\leq X$ we have
(39) $\mathbf{1}(x\in[0,\Delta^{-1}])\Big{(}1-e^{-\Delta^{a-2}}\Big{)}\leq|\mathcal{D}^{+}(x)|^{2}\leq\mathbf{1}(x\in[-\Delta^{-a},\Delta^{-1}+\Delta^{-a}])+e^{-\Delta^{a-2}},$
and
(40) $\mathbf{1}(x\in[\Delta^{-a},\Delta^{-1}-\Delta^{-a}])\Big{(}1-e^{-\Delta^{a-2}}\Big{)}\leq|\mathcal{D}^{-}(x)|^{2}\leq\mathbf{1}(x\in[0,\Delta^{-1}])+e^{-\Delta^{a-2}}.$
Proof.
We consider the entire functions $h^{\pm}$ from Lemma 1. We pick $b=2a$ and $c=3a$. Expressing $h^{\pm}$ in terms of their Fourier transform, we get
$h^{\pm}(x)=\int e^{2\pi{\rm i}\xi x}\widehat{h^{\pm}}(\xi){\rm d}\xi=\sum_{\ell\geq 0}\frac{(2\pi{\rm i}x)^{\ell}}{\ell!}\int\xi^{\ell}\widehat{h^{\pm}}{\rm d}\xi=\sum_{\ell<\nu}\frac{(2\pi{\rm i}x)^{\ell}}{\ell!}\int\xi^{\ell}\widehat{h^{\pm}}(\xi){\rm d}\xi+\mathcal{R}(x).$
The cutoff $\nu$ will need to be suitably picked. By Taylor’s theorem, the remainder term satisfies the bound
$|\mathcal{R}|\leq\frac{(2\pi X)^{\nu}}{\nu!}\Big{|}\int\xi^{\nu}\widehat{h^{\pm}}(\xi){\rm d}\xi\Big{|}\leq\frac{(2\pi X)^{\nu}}{\nu!}\Delta^{c(\nu+2)},$
where we used Property iv of Lemma 1. Using Stirling’s formula, we see that by picking
(41) $\nu=100X\Delta^{c},$
we have
(42) $|\mathcal{R}|\leq e^{-100X\Delta^{c}}.$
Thus far we have established the following decomposition
$h^{\pm}(x)=\mathcal{D}^{\pm}(x)+\mathcal{R}(x),$
where $\mathcal{D}^{\pm}$ is a polynomial of degree at most $100X\Delta^{c}$. The $\ell$-th coefficient of $\mathcal{D}^{\pm}$, $\ell<\nu$, satisfies the bound
$\left|\frac{(2\pi{\rm i})^{\ell}}{\ell!}\int\xi^{\ell}\widehat{h^{\pm}}(\xi){\rm d}\xi\right|\leq\frac{(2\pi)^{\ell}}{\ell!}\Delta^{c(\ell+2)},$
by Property iv of Lemma 1.
It remains to prove the inequalities (39) and (40). We prove the left-hand side inequality of (39), the other ones are proved the same way. The case $x\notin[0,\Delta^{-1}]$ is clear. When $x\in[0,\Delta^{-1}]$, we have from Property ii of Lemma 1 and (42) that
$|\mathcal{D}^{+}(x)|=|h^{+}(x)-\mathcal{R}(x)|\geq 1-e^{-\Delta^{2a-1}}-e^{-100X\Delta^{c}}.$
This is $1-e^{-\Delta^{a-2}}$ for $\Delta$ large enough. ∎
We now apply the last two lemmas to the proof of Proposition 2.
Proof of Proposition 2.
For simplicity, we prove the case $\ell=\mathcal{L}$. The proof when $\ell<\mathcal{L}$ is the same. Also, without loss of generality, we can assume $\alpha\geq 1$ by replacing $\alpha$ by $\alpha\vee 1$ in what follows.
Define the increments
[…]
For $\ell=0$, we take
(43) $Y_{0}=|\mathcal{M}_{0}|^{2}.$
The technical reason behind this is that $\mathcal{M}_{0}$ is a short Dirichlet polynomial, but $|\mathcal{M}_{0}^{-1}|$ and $\log|\mathcal{M}_{0}^{-1}|$ are not. Therefore, it is technically easier to approximate events involving $|\mathcal{M}_{0}|^{2}$.
Consider the events $G_{\ell}$ as in (21) and $G(V)$ as in (20). We want to express these in terms of the increments
[…]
The set $\mathcal{I}=\mathcal{I}(V)$ of tuples $\mathbf{u}=(u_{\ell},0\leq\ell\leq\mathcal{L})$ with $u_{\ell}\in\Delta_{\ell}^{-1}\cap\mathbb{Z}$ must be picked to mimic the restrictions in the $G_{\ell}$’s and $G(V)$. For $\ell=0$, from the definition of $G_{0}$ in (20), we take $u_{0}\in[L_{0},U_{0}]$. Clearly, $|u_{0}|\leq 1$. We discretize this interval with a step size of
(44) $\Delta_{0}^{-1}=e^{-100\alpha t_{0}}.$
Note that for $Y_{0}\in[u_{0},u_{0}+\Delta_{0}^{-1}]$, we then have by expanding the logarithm
$-\frac{1}{2}\log u_{0}-\frac{1}{2}\Delta_{0}^{-1/2}\leq-\frac{1}{2}\log Y_{0}\leq-\frac{1}{2}\log u_{0}$
For the $u_{\ell}$, $1\leq\ell\leq\mathcal{L}$, we ask that the $u$’s satisfy the “barrier” condition,
(45) $L_{k}(u_{0})\leq u_{0}+\sum_{1\leq k\leq\ell}u_{k}\leq U_{k}(u_{0})\text{ for $1\leq\ell\leq\mathcal{L}$.}$
In the following, to ease notation, we write $L_{\ell}(u_{0})=L_{\ell}$. The condition (45) implies the following bounds on the $u_{\ell}$’s, $1\leq\ell\leq\mathcal{L}$:
$u_{\ell}$ $\leq U_{\ell}-L_{\ell-1}=(\mathfrak{s}\,m(u_{0})+\mathcal{C})\log_{\ell-1}t-(\mathfrak{s}\,m(u_{0})-\mathcal{B})\log_{\ell}t,$
$u_{\ell}$ $\geq L_{\ell}-U_{\ell-1}=(\mathfrak{s}\,m(u_{0})-\mathcal{B})\log_{\ell-1}t-(\mathfrak{s}\,m(u_{0})+\mathcal{C})\log_{\ell}t.$
From the choice of slope $m(z)$ in (26) and $\mathcal{B}$ and $\mathcal{C}$ in (27), we can assume that
(46) $|u_{\ell}|\leq 2\alpha\mathfrak{s}\log_{\ell-1}t,\quad 1\leq\ell\leq\mathcal{L}.$
This is the important input for the choice of parameters. From this, we take
$\Delta_{\ell}=10\alpha\mathfrak{s}\log_{\ell-1}t,\quad 1\leq\ell\leq\mathcal{L}.$
This ensures that
(47) $\sum_{1\leq k\leq\ell}\Delta_{k}^{-1}\leq\alpha^{-1},\quad 1\leq\ell\leq\mathcal{L}.$
This choice is also motivated by the condition $|u_{\ell}\Delta_{\ell}^{-1}|\leq 1$ arising later in (59). Finally, the restriction $\{\mathcal{S}_{\mathcal{L}}+\log|\mathcal{M}_{0}^{-1}|>V\}$ imposes the following condition on $\mathcal{I}$
(48) $\sum_{1\leq\ell\leq\mathcal{L}}u_{\ell}-\frac{1}{2}\log u_{0}>V\ .$
With these choices, it is not hard to check that we have the following inclusions:
$\bigcup_{\mathbf{u}\in\mathcal{I}}\{Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}],\ell\leq\mathcal{L}\}\subset$
(49) $\{S_{\mathcal{L}}-\frac{1}{2}\log Y_{0}>V-\alpha^{-1},S_{\ell}\in[L_{\ell},U_{\ell}+\alpha^{-1}],1\leq\ell\leq\mathcal{L},Y_{0}\in[L_{0},U_{0}+\Delta_{0}^{-1}]\}\ ,$
and
$\bigcup_{\mathbf{u}\in\mathcal{I}}\{Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}],\ell\leq\mathcal{L}\}\supset$
(50) $\{S_{\mathcal{L}}-\frac{1}{2}\log Y_{0}>V+\alpha^{-1},S_{\ell}\in[L_{\ell}+\alpha^{-1},U_{\ell}],1\leq\ell\leq\mathcal{L},Y_{0}\in[L_{0}+\Delta_{0}^{-1},U_{0}]\}\ .$
Note that the events on the left are disjoint for two different tuples $\mathbf{u}$.
The lower and upper bound in the statement of Proposition 2 are proved separately using the two above inclusions. Both proofs consist in re-expressing the probabilities of the increments $Y_{k}$ in terms of the ones of a random model where we effectively replace $p^{-i\tau}$ by a uniform random variable on the unit circle. This will be achieved by converting the indicator functions in terms of polynomials with Lemma 2 and using mean-value theorems for Dirichlet polynomials. The random model is defined as follows: for $(\theta_{p},p\text{ primes})$ be IID random variables uniformly distributed on $[0,2\pi]$, we consider the random model for the Dirichlet sums $S_{\ell}$
(51) $\mathcal{S}_{\ell}=\sum_{T_{0}<p\leq\exp(e^{t_{\ell}})}\frac{{\rm Re}\ e^{{\rm i}\theta_{p}}}{p^{\sigma}}+\frac{{\rm Re}\ e^{2{\rm i}\theta_{p}}}{2p^{2\sigma}},\quad 1\leq\ell\leq\mathcal{L},$
and their corresponding increments
[…]
For $\ell=0$, keeping in line with the choice in (43), we define
$\mathcal{Y}_{0}=\prod_{p\leq T_{0}}\Big{|}1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\Big{|}^{2}.$
Proof of $\mathbf{P}(G(V))\gg\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V)$.
It suffices to prove that
(53) $\mathbf{P}(G(V))\gg\mathbf{P}(\mathcal{G}(V+\alpha^{-1}))$
for the event $\mathcal{G}(V)$ given in (23). Indeed, note that the added terms $+\alpha^{-1}$ to the barrier $U_{\ell}$ and $L_{\ell}$ are irrelevant by Proposition 1. Moreover, the same proposition implies that $\mathbf{P}(\mathcal{G}(V+\alpha^{-1}))\gg\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V+\alpha^{-1})$. This is $\gg\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V)$ by (18).
To prove (53), we use the inclusion (49) to get
$\mathbf{P}(G(V))\gg\sum_{\mathbf{u}\in\mathcal{I}}\mathbf{P}(Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}],0\leq\ell\leq\mathcal{L}).$
To apply Lemma 2 to the indicator function $[0,\Delta_{\ell}^{-1}]$, we introduce the restriction $|Y_{\ell}-u_{\ell}|\leq X_{\ell}$ for a parameter $X_{\ell}\leq 100\Delta_{\ell}^{4a}$. The justification of the constraint will be specified in (72). We will also need to pick $a=5$, to guarantee that $\Delta_{\ell}^{a-2}>\Delta_{\ell}^{2}$. Together with (55), this informs the choice (12). With this, we have for $1\leq\ell\leq\mathcal{L}$
$\mathbf{1}(Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}])$ $\geq\mathbf{1}_{[0,\Delta_{\ell}^{-1}]}(Y_{\ell}-u_{\ell})\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})$
$\geq|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell}),$
where $\mathcal{D}_{\ell}^{-}(x)$ is a polynomial of degree $\leq 100X_{\ell}\Delta_{\ell}^{3a}$ as guaranteed by Lemma 2. For $\ell=0$, there is no need to introduce the indicator function. Indeed, we have the pointwise bound
$|Y_{0}|=\prod_{p\leq T_{0}}\Big{|}1-\frac{p^{-{\rm i}\tau}}{p^{\sigma}}\Big{|}^{2}\leq\exp(10T_{0}^{1/2}).$
Therefore, by picking $X_{0}=\exp(10T_{0}^{1/2})=\exp(10\sqrt{\log\log T})$, we get
$\mathbf{1}(Y_{0}\in[u_{0},u_{0}+\Delta_{0}^{-1}])\geq|\mathcal{D}_{0}^{-}(Y_{0}-u_{0})-e^{-\Delta_{0}^{a-2}}|^{2}.$
This means that $\prod_{0\leq\ell\leq\mathcal{L}}\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})$ is a Dirichlet polynomial of length less than
(54) $\exp(2\cdot 100e^{t_{\mathcal{L}}}X_{\mathcal{L}}\Delta_{\mathcal{L}}^{3a})=\exp\Big{(}2\cdot 10^{4}\frac{(\alpha\vee 1)^{7a}(\mathfrak{s}\log_{\mathcal{L}}t)^{7a}}{e^{\mathfrak{s}\log_{\mathcal{L}}t}}\Big{)}.$
To ensure that the length is less than $T^{1/100}$, we are led to impose
(55) $\boxed{\mathfrak{s}\log_{\mathcal{L}}t-7a\log(\mathfrak{s}\log_{\mathcal{L}}t)>7a\log(\alpha\vee 1)+\log(2\cdot 10^{6}).}$
For $\alpha$ large, this essentially imposes $\mathfrak{s}\log_{\mathcal{L}}t$ larger than $7a\log(\alpha\vee 1)$. Our goal is now to get rid of the indicator function for $1\leq\ell\leq\mathcal{L}$. In this case, writing $\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})=1-\mathbf{1}(|Y_{\ell}-u_{\ell}|>X_{\ell})$, we get by expanding the product
$\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{1}(|Y_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
(56) $=\sum_{J\subseteq\{1,\dots,\mathcal{L}\}}(-1)^{|J|}\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{j\in J}\mathbf{1}(|Y_{j}-u_{j}|>X_{j})\right].$
The dominant term will be the summand corresponding to $J=\emptyset$. It is good to start with estimating this term. By the mean-value theorem of Dirichlet polynomials, see Lemma 7, the first term can be expressed in terms of the random increments $\mathcal{Y}_{\ell}$, which are now independent:
$\mathbf{E}\left[\prod_{\ell}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]=(1+\operatorname{O}(T^{-1/100}))\prod_{\ell}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right].$
Each factor in the product is easily bounded below by
(57) $\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]\geq\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{(}1-\frac{2e^{-\Delta_{\ell}^{a-2}}}{[\mathbf{E}[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}]]^{1/2}}\Big{)}.$
We can apply Lemma 2 again to convert the polynomial back to an indicator function
$\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$ $\geq\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\mathbf{1}(|\mathcal{Y}_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
$\geq(1-e^{-\Delta_{\ell}^{a-2}})\mathbf{P}(\mathcal{Y}_{\ell}-u_{\ell}\in[\Delta_{\ell}^{-a},\Delta_{\ell}^{-1}-\Delta_{\ell}^{-a}]),$
where we noticed that $\mathcal{Y}_{\ell}-u_{\ell}\in[\Delta_{\ell}^{-a},\Delta_{\ell}^{-1}-\Delta_{\ell}^{-a}]$ implies $|\mathcal{Y}_{\ell}-u_{\ell}|\leq X_{\ell}$. The goal is to rewrite the probability in terms of Gaussian increments and without the overspill $\Delta_{\ell}^{-a}$. This is evaluated differently for $\ell=0$, $\ell=1$ and $1<\ell\leq\mathcal{L}$.
For the range $1<\ell\leq\mathcal{L}$, by Lemma 8, the probability is
(58) $\mathbf{P}(\mathcal{N}_{\ell}-u_{\ell}\in[\Delta_{\ell}^{-a},\Delta_{\ell}^{-1}-\Delta_{\ell}^{-a}])+e^{-ce^{t_{\ell}/2}},$
for the Gaussian increment $\mathcal{N}_{\ell}=\mathcal{Z}_{\ell}-\mathcal{Z}_{\ell-1}$. To get rid of the overspills $\Delta_{\ell}^{-a}$, note that by the choice of $u_{\ell}$ and $\Delta_{\ell}$, we have
(59) $|u_{\ell}|\Delta_{\ell}^{-a}\leq 1,$
so that
(60) $\mathbf{P}(\mathcal{N}_{\ell}-u_{\ell}\in[\Delta_{\ell}^{-a},\Delta_{\ell}^{-1}-\Delta_{\ell}^{-a}])=(1+\operatorname{O}(\Delta_{\ell}^{-a}))\mathbf{P}(\mathcal{N}_{\ell}-u_{\ell}\in[0,\Delta_{\ell}^{-1}]).$
Moreover, we have trivially
(61) $\mathbf{P}(\mathcal{N}_{\ell}-u_{\ell}\in[0,\Delta_{\ell}^{-1}])\gg\Delta_{\ell}^{-1}e^{-u_{\ell}^{2}/2v_{\ell}}\gg e^{-\Delta_{\ell}^{2}},$
by the choice of $\Delta_{\ell}$. This is much larger than the additive error in (58). We conclude that for each $1<\ell\leq\mathcal{L}$
(62) $\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\geq(1+\operatorname{O}(\Delta_{\ell}^{-a}))\mathbf{P}(\mathcal{N}_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}])\gg e^{-\Delta_{\ell}^{2}}.$
Putting this together with the choice $a=5$ allows us to rewrite (57) as
$\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]$ $\geq\Big{(}1-e^{-\Delta_{\ell}^{a-2}/2}\Big{)}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$
(63) $\geq(1+\operatorname{O}(\Delta_{\ell}^{-a}))\mathbf{P}(\mathcal{N}_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}]).$
For $\ell=1$, the additive error in Lemma 8 is too big compared to the probability. Taking a larger $t_{0}$ to reduce the error would cause the polynomial $\mathcal{M}_{0}$ to be too long. To circumvent this issue, we estimate the probability using Lemma 9. This is done at the cost of a multiplicative constant, which is harmless when it appears for a single term in the product over $\ell$. We ultimately get
(64) $\mathbf{E}\left[|\mathcal{D}_{1}^{-}(\mathcal{Y}_{1}-u_{1})-e^{-\Delta_{1}^{a-2}}|^{2}\right]\gg\mathbf{P}(\mathcal{N}_{1}\in[u_{1},u_{1}+\Delta_{1}^{-1}]).$
We now establish a similar bound for $\mathcal{Y}_{0}$. Lemma 2 implies that
$\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})|^{2}\right]\gg\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-a},u_{0}+\Delta_{0}^{-1}-\Delta_{0}^{-a}])\geq\frac{1}{2}\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}]).$
To bound the probability from below, note that
$\mathbf{E}[\mathcal{Y}_{0}]\geq 1,\qquad\mathbf{E}[\mathcal{Y}_{0}^{2}]=\mathbf{E}[|\mathcal{M}_{0}|^{4}]=\prod_{p\leq T_{0}}\Big{(}1+\frac{2}{p^{2\sigma}}+\operatorname{O}(p^{-3\sigma})\Big{)}\ll\exp(2t_{0}).$
This implies
(65) $\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}])\gg\Delta_{0}^{-1}\mathbf{P}(\mathcal{Y}_{0}>u_{0}+2\Delta_{0}^{-1})\gg\Delta_{0}^{-1}\frac{\mathbf{E}[\mathcal{Y}_{0}]^{2}}{\mathbf{E}[\mathcal{Y}_{0}^{2}]}\gg e^{-\Delta_{0}^{a-2}}.$
This ensures that in (57)
(66) $\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})-e^{-\Delta_{0}^{a-2}}|^{2}\right]\gg\mathbf{E}\left[|\mathcal{D}_{0}^{-}(\mathcal{Y}_{0}-u_{0})|^{2}\right]\gg\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}]).$
Altogether, Equations (63), (64) and (66) show that the term for $J=\emptyset$ in (56) is
(67) $\gg\mathbf{P}(\mathcal{Y}_{0}\in[u_{0}+\Delta_{0}^{-1},u_{0}+2\Delta_{0}^{-1}])\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{P}(\mathcal{N}_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}]).$
It remains to estimate the summands in (56) with $J\neq\emptyset$. For each $J$, we can bound the indicator functions to get
(68) $\leq\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(Y_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{j\in J}X_{j}^{-2q_{j}}|Y_{j}-u_{j}|^{2q_{j}}\right],$
for an appropriate choice of $q_{j}$, $j\in J$. The length of the polynomial in the expectation is then at most
(69) $\exp\left(\sum_{0\leq\ell\leq\mathcal{L}}2\cdot 10^{2}e^{t_{\ell}}X_{\ell}\Delta_{\ell}^{3a}+\sum_{j\leq\mathcal{L}}2e^{t_{j}}(2q_{j})\right).$
The mean-value theorem of Dirichlet polynomial can be applied if $q_{j}$ is not too large, for example $q_{j}\leq X_{j}\Delta_{j}$ would do. Passing to the random model, Equation (68) is
(70) $=\prod_{\ell\in J^{c}\cup\{0\}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\right]\prod_{j\in J}X^{-2q_{j}}\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{2}|\mathcal{Y}_{j}-u_{j}|^{2q_{j}}\right].$
The expectation on the right is for each $j$
(71) $\leq\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{4}\right]^{1/2}\cdot\mathbf{E}[|\mathcal{Y}_{j}-u_{j}|^{4q_{j}}]^{1/2}.$
We evaluate both terms separately. By Equations (62) for $j>1$ and Equations (64), (65) for $j=1$ and $j=0$, we have
$\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}\right]^{1/2}\geq\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{2}\right]\gg e^{-\Delta_{j}^{2}}.$
This implies
$\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})-e^{-\Delta_{j}^{a-2}}|^{4}\right]^{1/2}$ $\leq 2\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}\right]^{1/2}+2e^{-2\Delta_{j}^{a-2}}$
$\leq 4\mathbf{E}[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}]^{1/2}.$
The fourth moment for $j\geq 1$ is estimated using the bounds on the coefficients in Lemma 2 and the elementary estimates on the random model, see Equation (113) in Lemma 10,
$\mathbf{E}[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{4}]$ $\leq\Delta_{j}^{4a}\mathbf{E}\left[\Big{(}\sum_{\ell\geq 0}\frac{(2\pi\Delta_{j}^{2a})^{\ell}}{\ell!}|\mathcal{Y}_{j}-u_{j}|^{\ell}\Big{)}^{4}\right]$
$\leq\Delta_{j}^{4a}\mathbf{E}[\exp(8\pi\Delta_{j}^{2a}|\mathcal{Y}_{j}-u_{j}|)]$
$\leq\Delta_{j}^{4a}e^{8\pi\Delta_{j}^{2a}|u_{j}|+8^{2}\pi^{2}\Delta_{j}^{4a}v_{j}^{2}/2}$
$\leq\exp(100\Delta_{j}^{4a}),$
since $|u_{j}|\leq\Delta_{j}$. For the second factor in (71), Minkowski’s inequality implies
$\mathbf{E}[|\mathcal{Y}_{j}-u_{j}|^{4q_{j}}]^{1/(4q_{j})}\leq\mathbf{E}[|\mathcal{Y}_{j}|^{4q_{j}}]^{1/(4q_{j})}+|u_{j}|.$
Moreover, the bound (111) yields $\mathbf{E}[|\mathcal{Y}_{j}|^{4q_{j}}]\leq\frac{(4q_{j})!}{2^{2q_{j}}(2q_{j})!}v_{j}^{4q_{j}}$, which by Stirling gives the estimate
$\mathbf{E}[|\mathcal{Y}_{j}|^{4q_{j}}]^{1/(4q_{j})}\leq q_{j}^{1/2}.$
We are now in position to pick $q_{j}=X_{j}$. Note that this satisfies the restriction $q_{j}\leq X_{j}\Delta_{j}^{3a}$ imposed by (69). With this choice, we have
$X^{-4q_{j}}\mathbf{E}[|\mathcal{Y}_{j}-u_{j}|^{4q_{j}}]\leq\Big{(}\frac{1}{X_{j}^{1/2}}+\frac{|u_{j}|}{X_{j}}\Big{)}^{4q_{j}}\leq e^{-4X_{j}}.$
Putting all these observations back in Equation (70) gives the upper bound
$\leq\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\prod_{j\in J}\frac{\exp(50\Delta_{j}^{4a}-2X_{j})}{\mathbf{E}\left[|\mathcal{D}_{j}^{-}(\mathcal{Y}_{j}-u_{j})|^{2}\right]}.$
This leads to the choice
(72) $X_{j}=100\Delta_{j}^{4a}.$
Together with the bound (62), this shows that the above is $\leq\prod_{j\in J}e^{-\Delta_{j}^{4a}}$. Combining this with the dominant term (63), we obtain
$\mathbf{E}\left[\prod_{0\leq\ell\leq\mathcal{L}}|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})-e^{-\Delta_{\ell}^{a-2}}|^{2}\prod_{1\leq\ell\leq\mathcal{L}}\mathbf{1}(|\mathcal{Y}_{\ell}-u_{\ell}|\leq X_{\ell})\right]$
$\geq\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{)}\Big{(}1-e^{-\Delta^{a}_{\mathcal{L}}/2}-\sum_{J\subset\{1,\dots,\mathcal{L}\},J\neq\emptyset}\prod_{j\in J}e^{-\Delta_{j}^{4a}}\Big{)}$
$=\Big{(}\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]\Big{)}\Big{(}2-e^{-\Delta_{\mathcal{L}}^{a}/2}-\prod_{j\leq\mathcal{L}}(1+e^{-\Delta_{j}^{4a}})\Big{)}.$
This is $\gg\prod_{0\leq\ell\leq\mathcal{L}}\mathbf{E}\left[|\mathcal{D}_{\ell}^{-}(\mathcal{Y}_{\ell}-u_{\ell})|^{2}\right]$ for $\Delta_{\mathcal{L}}$ large enough. The final result is Equation (67). To re-express the events in terms of the random model, we use the inclusion (49) with $\mathcal{Y}_{0}$ and $\mathcal{Z}_{\ell}$. This proves the claim (53).
Proof of $\mathbf{P}(G(V))\ll\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V)$. The proof of the upper bound follows the same method using the polynomial $\mathcal{D}^{+}$ instead of $\mathcal{D}^{-}$. The proof is in fact slightly easier as we do not need to bound a long polynomial as in (68). The corresponding quantity needs to be estimated in the random model where the factors are decoupled. ∎
We will need the following spinoff of Proposition 1 in the proof of Proposition 4.
Corollary 2.
Consider the complex partial sums $\widetilde{S}_{k}=\sum_{T_{0}<p\leq T_{k}}\frac{p^{-{\rm i}\tau}}{p^{\sigma}}+\frac{p^{-2{\rm i}\tau}}{2p^{2\sigma}}$, $1\leq\ell\leq\mathcal{L}$. For $1<\ell\leq\mathcal{L}$ and $q\leq 100\alpha^{2}(t_{\ell}-t_{\ell-1})$, we have
$\mathbf{E}[|\widetilde{S}_{\ell}-\widetilde{S}_{\ell-1}|^{2q}\mathbf{1}(G_{\ell-1})]\ll\mathbf{E}[|\widetilde{S}_{\ell}-\widetilde{S}_{\ell-1}|^{2q}]\mathbf{P}(\{\mathcal{Z}_{\ell-1}\geq L_{\ell-1}-(\alpha\vee 1)^{-1}\}\cap\mathcal{G}_{0}),$
Proof.
This is done the same way as the proof of Proposition 2 by approximating the indicator function $\mathbf{1}(G_{\ell-1})$ by polynomials. The mean-value theorem for Dirichlet polynomial can be applied with the additional $(\widetilde{S}_{\ell}-\widetilde{S}_{\ell-1})^{q}$ since the total length is still admissible, as in Equations (54) and (69), by the bound on $q$. Since we only need an upper bound, one can ultimately drop all restrictions on $\mathcal{Z}_{\ell}$, $1\leq\ell\leq\mathcal{L}$, except for the lower bound on $\mathcal{Z}_{\ell-1}$. The restriction on $\mathcal{Z}_{0}$ also remains. ∎
3.3. Proof of Proposition 3
The main tool to prove the proposition is Lemma 5 from the appendix. The following lemma adapts it to our setting.
Lemma 3.
Consider the mollifier $\mathcal{M}$ defined in (31). For $\alpha>0$, let $G(V)$ be the event in (23) for $V\sim\alpha\log\log T$ and $\mathcal{Z}_{\ell}$, $0\leq\ell\leq\mathcal{L}$, the random variables defined in (15) and (14). We have for $\delta>\delta^{\star}(\alpha)$ given in (33)
$\mathbf{E}\left[\Big{|}\zeta\mathcal{M}-1\Big{|}^{2}\mathbf{1}(G(V))\right]\ll\left(\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\right)\cdot\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V).$
Proof.
The result follows from Lemma 5 after properly approximating the indicator function. To do this, we proceed as in the proof of Proposition 2 by writing $\mathbf{1}(G_{\mathcal{L}})$ as a Dirichlet polynomial. Writing the event in terms of the increments $Y_{\ell}$ and using the inclusion (49) yields
$\sum_{\mathbf{u}\in\mathcal{I}}\mathbf{E}\left[\Big{|}\zeta\mathcal{M}-1\Big{|}^{2}\mathbf{1}(Y_{\ell}\in[u_{\ell},u_{\ell}+\Delta_{\ell}^{-1}],\ell\leq\mathcal{L})\right]$
$\hskip 56.9055pt\leq 2\sum_{\mathbf{u}\in\mathcal{I}}\mathbf{E}\left[\Big{|}\zeta\mathcal{M}-1\Big{|}^{2}\prod_{\ell\leq\mathcal{L}}|\mathcal{D}_{+}(Y_{\ell}-u_{\ell})|^{2}\right]$
where the inequality follows by an application of Lemma 2. In view of the choice of $\mathfrak{s}$ in (55), the product is a well-factorable polynomial and of length admissible to apply Lemma 5, cf. Equation (54). The above is then
$\ll\Big{(}\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\Big{)}\sum_{\mathbf{u}\in\mathcal{I}}\mathbf{E}\left[\prod_{\ell\leq\mathcal{L}}|\mathcal{D}_{+}(Y_{\ell}-u_{\ell})|^{2}\right].$
We can also replace $Y_{\ell}$ by $\mathcal{Y}_{\ell}$ at a cost of a small multiplication constant $1+\operatorname{O}(T^{-99})$. It remains to re-express the polynomial in terms of the $\mathcal{Y}_{\ell}$’s as in the proof of Proposition 2. This proves the claim. ∎
Proof of Proposition 3.
By introducing the event that $\zeta\mathcal{M}$ is close to $1$, we have
$\mathbf{P}(\log|\zeta|>V)$ $\geq\mathbf{P}\Big{(}\{\log|\mathcal{M}^{-1}|>V-\log|\zeta\mathcal{M}|\}\cap\{|\zeta\mathcal{M}-1|<\varepsilon\Big{)}$
$\geq\mathbf{P}\Big{(}\{\log|\mathcal{M}^{-1}|>V-\log(1-\varepsilon)\}\cap\{|\zeta\mathcal{M}-1|<\varepsilon\}\Big{)}$
If we also intersect with $G(V)$, we get the lower bound
$\geq\mathbf{P}\Big{(}\{\log|\mathcal{M}^{-1}|>V-\log(1-\varepsilon)\}\cap G(V)\Big{)}-\mathbf{P}\Big{(}\{|\zeta\mathcal{M}-1|>\varepsilon\}\cap G(V)\Big{)}$
The second term is bounded by Markov’s inequality
$\mathbf{P}\Big{(}\{|\zeta\mathcal{M}-1|>\varepsilon\}\cap G(V)\Big{)}\leq\frac{1}{\varepsilon^{2}}\mathbf{E}\Big{[}|\zeta\mathcal{M}-1|^{2}\mathbf{1}(G(V))\Big{]}$
The claimed bound then follows from Lemma 3. ∎
3.4. Proof of Proposition 4
The first step consists in applying Lemma 6. For this, it is necessary to consider the complex partial sums
(73) $\widetilde{S}_{\ell}=\sum_{T_{0}<p\leq T_{\ell}}\frac{p^{-{\rm i}\tau}}{p^{\sigma}}+\frac{p^{-2{\rm i}\tau}}{2p^{2\sigma}},\quad\ell\geq 0.$
We consider the decreasing events
[…]
with $A_{0}$ being the whole sample space. On $A_{\mathcal{L}}$, Lemma 6 applied with $A=10(\alpha\vee 1)$ implies
$\prod_{1\leq\ell\leq\mathcal{L}}|\mathcal{M}_{\ell}^{-1}|\geq\exp\Big{(}S_{\mathcal{L}}+\operatorname{O}^{\star}(5T_{0}^{-1/2})-\sum_{\ell=1}^{\mathcal{L}}e^{-10(\alpha\vee 1)(t_{\ell}-t_{\ell-1})}\Big{)}\geq\exp(S_{\mathcal{L}}-(\alpha\vee 1)^{-1}),$
where we use the notation of Lemma 6, writing $\operatorname{O}^{\star}$ for an error term with implicit constant smaller than $1$. These observations together means that
$\mathbf{P}\left(\{\log|\mathcal{M}^{-1}|>V\}\cap G(V)\right)$ $\geq\mathbf{P}\left(S_{\mathcal{L}}+\log|\mathcal{M}_{0}|^{-1}>V+(\alpha\vee 1)^{-1}\}\cap G(V)\right)-\mathbf{P}\left(G_{\mathcal{L}}\cap A_{\mathcal{L}}^{c}\right)$
$\geq\mathbf{P}\left(G(V+(\alpha\vee 1)^{-1})\right)-\mathbf{P}\left(G_{\mathcal{L}}\cap A_{\mathcal{L}}^{c}\right)$
$\gg\mathbf{P}\left(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\right)-\mathbf{P}\left(G_{\mathcal{L}}\cap A_{\mathcal{L}}^{c}\right).$
The second inequality is by definition of $G(V)$ and the third is by Propositions 1 and 2. It remains to bound the second probability. It is smaller than
$\sum_{\ell=1}^{\mathcal{L}}\mathbf{P}\Big{(}G_{\ell-1}\cap\{|\widetilde{S}_{\ell}-\widetilde{S}_{{\ell-1}}|>10\alpha(t_{\ell}-t_{\ell-1})\}\Big{)},$
with the convention that $G_{0}$ is the whole sample space $[T,2T]$. Markov’s inequality for fixed $q\in\mathbb{N}$ can be applied to each summand to yield
$\leq\frac{\mathbf{E}[|\widetilde{S}_{\ell}-\widetilde{S}_{{\ell-1}}|^{2q}\mathbf{1}(G_{\ell-1})]}{|10\alpha(t_{\ell}-t_{\ell-1})|^{2q}}\ll\frac{\mathbf{E}[|\widetilde{S}_{\ell}-\widetilde{S}_{{\ell-1}}|^{2q}]}{|10\alpha(t_{\ell}-t_{\ell-1})|^{2q}}\cdot\mathbf{P}\big{(}\mathcal{Z}_{\ell-1}\geq L_{\ell-1}-(\alpha\vee 1)^{-1},\mathcal{Z}_{0}>\alpha t_{0}\big{)},$
for $\ell>1$ by Corollary 2. The factor $(\alpha\vee 1)^{-1}$ is irrelevant by (18). For $\ell=1$, we simply use the bound $\frac{\mathbf{E}[|\widetilde{S}_{\ell}-\widetilde{S}_{{\ell-1}}|^{2q}]}{|10\alpha(t_{\ell}-t_{\ell-1})|^{2q}}$. The $2q$-moment can be estimated using Lemma 11. Picking $q=\lfloor\alpha^{2}(t_{\ell}-t_{\ell-1})\rfloor$, this yields the upper bound
$\ll\alpha\sqrt{t_{\ell}-t_{\ell-1}}\exp\Big{(}\frac{-100\alpha^{2}(t_{\ell}-t_{\ell-1})^{2}}{t_{\ell}-t_{\ell-1}}\Big{)}\mathbf{P}(\mathcal{Z}_{\ell-1}\geq L_{\ell-1},\mathcal{Z}_{0}>\alpha t_{0})$
$\ll e^{-99\alpha^{2}(t_{\ell}-t_{\ell-1})}\frac{e^{-2\alpha^{2}\delta}}{\alpha\sqrt{t}}\exp(-\alpha^{2}(t_{\ell-1}-t_{0})+2\alpha\mathcal{C}\log_{\ell-1}t)\mathbf{P}(\mathcal{Z}_{0}>\alpha t_{0})$
$\ll\frac{e^{-2\alpha^{2}\delta}e^{-\alpha^{2}t}}{\alpha\sqrt{t}}e^{-10\alpha^{2}\mathfrak{s}\log_{\ell-1}t}\cdot e^{\alpha^{2}t_{0}}\mathbf{P}(\mathcal{Z}_{0}>\alpha t_{0}).$
The second inequality follows by a Gaussian estimate and the variance estimate 17. The factor $e^{\alpha^{2}t_{0}}\mathbf{P}(\mathcal{Z}_{0}>\alpha t_{0})$ is $\ll a_{\alpha}e^{\gamma\alpha^{2}}$ by (116). Summing the above bound on $1\leq\ell\leq\mathcal{L}$ yields
$\mathbf{P}\left(G_{\mathcal{L}}\cap A_{\mathcal{L}}^{c}\right)\ll\frac{C_{\alpha}(\delta)}{\alpha\sqrt{t}}e^{-\alpha^{2}t}e^{-10\alpha^{2}\mathfrak{s}\log_{\mathcal{L}}t}$
By the choice of $\mathfrak{s}$, we thus have
$\mathbf{P}\left(\{\log|\mathcal{M}^{-1}|>V\}\cap G(V)\right)\gg\mathbf{P}\left(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>V\right)-\mathbf{P}\left(G_{\mathcal{L}}\cap A_{\mathcal{L}}^{c}\right)\gg C_{\alpha}(\delta)\frac{e^{-\alpha^{2}t}}{\alpha\sqrt{t}},$
as claimed
4. Proof of Corollary 1
We first establish that, for $k\geq 0$, $\sigma=1/2+\delta e^{-t_{\mathcal{L}}}$, and $\delta$ satisfying the conditions of Theorem 1,
(74) $\mathbf{E}[|\zeta(\sigma+{\rm i}\tau)|^{2k}]\geq c_{k}(\log T)^{k^{2}},$
with the coefficient proportional to
(75) $a_{k}e^{\gamma k^{2}}(k+1)^{-38k^{2}}.$
Firstly, we have
[…]
$\geq 2k\int_{\alpha^{-}}^{\alpha^{+}}e^{2k\alpha t}\mathbf{P}(\log|\zeta(\sigma+{\rm i}\tau)|\geq\alpha t)t\;{\rm d}\alpha$
$\gg 2k\int_{\alpha^{-}}^{\alpha^{+}}\frac{C_{\alpha}(\delta)}{\alpha}e^{2k\alpha t}e^{-\alpha^{2}t}\sqrt{t}\;{\rm d}\alpha,$
by localising to a window around $\alpha t$ and employing Theorem 1. The values $\alpha^{-},\alpha^{+}$ can be taken to be
$\alpha^{-}=k-\frac{1}{\sqrt{t}},\qquad\alpha^{+}=k.$
For $\alpha>0$, $C_{\alpha}(\delta)$ is decreasing and therefore the above is greater than
$\gg C_{k}(\delta)e^{k^{2}t}\int_{\alpha^{-}}^{\alpha^{+}}e^{-(\alpha-k)^{2}t}\sqrt{t}\;{\rm d}\alpha\gg C_{k}(\delta)e^{k^{2}t}.$
As noted in (19) the coefficient grows as (75) using the optimal $\delta\sim\log(\alpha+1)$, completing the proof of (74).
The corollary follows by demonstrating that the moments on-axis are an upper bound for the moments off-axis. To do so, we employ a result from [AOR21], a generalization of a convexity result of Gabriel [Gab27].
Proposition 5 (Corollary 3.4 of [AOR21]).
Write $z=v+it$. Suppose that $F$ is a complex valued function, analytic in the half-plane $v\geq 1/2$, such that $|F(z)|\rightarrow 0$ uniformly in $v>1/2$ as $t\rightarrow\infty$ and
$\int_{\mathbb{R}}|F(v+{\rm i}t)|^{k}{\rm d}t\rightarrow 0$
as $v\rightarrow\infty$, for $k>0$. Then for any $v\geq 1/2$ and $k>0$
$\int_{\mathbb{R}}|F(v+{\rm i}t)|^{k}{\rm d}t\leq\int_{\mathbb{R}}|F(1/2+{\rm i}t)|^{k}{\rm d}t.$
The purpose of the following lemma, an adaptation of Lemma 3.5 of [AOR21], is to construct an analytic approximation for the indicator function of the rectangle
$\{\nu+{\rm i}t:\nu\in[\tfrac{1}{2},\sigma],t\in[T,2T]\}.$
Lemma 4 (Adapted from [AOR21]).
Let $b_{1}\in(0,1)$ and $\Delta,T,A,b_{2}>0$. There exists an entire function $\Phi_{\Delta,T}(z)\equiv\Phi(z)$ such that, for $z=\nu+{\rm i}t$ with $\nu\geq\frac{1}{2}$ and $t\in\mathbb{R}$,
  1. (i)
    For $|t-\frac{3}{2}T|\geq\frac{3}{4}(1+b_{2})T$, uniformly in $\nu\geq\frac{1}{2}$, $\Phi(z)\ll_{A}b_{2}^{-A}\Delta^{1-A}$.
  2. (ii)
    For $|t-\frac{3}{2}T|\leq\frac{3}{4}(1-b_{1})T$, $|\Phi(z)|=1+\operatorname{O}_{b_{1},A}(\Delta^{-A})+\operatorname{O}((\nu-\tfrac{1}{2})\frac{\Delta^{2}}{T})$.
  3. (iii)
    For $|t-\frac{3}{2}T|\leq\frac{3}{4}(1+b_{2})T$, $|\Phi(z)|\ll 1+(\nu-\frac{1}{2})\frac{\Delta^{2}}{T}$.
  4. (iv)
    $\Phi(z)\rightarrow 0$ uniformly in $t$ as $\nu\rightarrow\infty$.
In order to apply Proposition 5, we use a Dirichlet polynomial approximation of $\zeta(\sigma+{\rm i}t)$. Define for $\sigma\geq\frac{1}{2}$, $t\asymp T$
$D(\sigma+{\rm i}t)=\begin{cases}\sum_{n\leq T^{2}}n^{-\sigma-{\rm i}t}w\left(\log\frac{e^{100}n}{T^{2}}\right),&\frac{1}{2}\leq\sigma\leq 2\\ \sum_{n\leq T}n^{-\sigma-{\rm i}t},&\sigma>\frac{1}{2},\end{cases}$
where the smoothing $w(x)$ is
$w(x)=\frac{1}{2\pi{\rm i}}\int_{2-{\rm i}\infty}^{2+{\rm i}\infty}e^{-xz}\left(\frac{e^{z}-1}{z}\right)^{100}\frac{{\rm d}z}{z}.$
By Lemma 2.8 of [AOR21], we then have
(76) $\zeta(\sigma+{\rm i}t)=D(\sigma+{\rm i}t)+\operatorname{O}\left(\tfrac{1}{T}\right).$
Set $F(z)=D(z)\Phi(z)$. By properties (i) and (iv) of Lemma 4, the requirements of Proposition 5 are satisfied. Thus for any $k>0$ we have the ‘smoothed’ version of the required bound:
(77) $\int_{\mathbb{R}}|D(\sigma+{\rm i}t)\Phi_{\Delta,T}(\sigma+{\rm i}t)|^{2k}{\rm d}t\leq\int_{\mathbb{R}}|D(1/2+{\rm i}t)\Phi_{\Delta,T}(1/2+{\rm i}t)|^{2k}{\rm d}t.$
The aim is now to unsmooth both sides using Lemma 4.
By property (ii), with $\Delta\asymp T\log T$ and $b_{1}=\frac{1}{3}$, and the approximation (76), we get the lower bound
(78) $\int_{T}^{2T}|\zeta(\sigma+{\rm i}t)|^{2k}{\rm d}t\ll\int_{T}^{2T}|D(\sigma+{\rm i}t)|^{2k}{\rm d}t\ll\int_{\mathbb{R}}|D(\sigma+{\rm i}t)\Phi_{\Delta,T}(\sigma+{\rm i}t)|^{2k}{\rm d}t.$
The upper bound follows similarly, using properties (i), (iii), and (iv) of Lemma 4,
$\int_{\mathbb{R}}|D(1/2+{\rm i}t)\Phi_{\Delta,T}(1/2+{\rm i}t)|^{2k}{\rm d}t$ $\ll_{A}\Delta^{2k(1-A)}b_{2}^{-2kA}\int_{|t-\frac{3}{2}T|>\frac{3}{4}T(1+b_{2})}|D(1/2+{\rm i}t)|^{2k}{\rm d}t$
$\qquad+\int_{|t-\frac{3}{2}T|\leq\frac{3}{4}T(1+b_{2})}|D(1/2+{\rm i}t)|^{2k}{\rm d}t$
$\ll\int_{|t-\frac{3}{2}T|\leq T}|D(1/2+{\rm i}t)|^{2k}{\rm d}t$
$\ll\int_{T}^{2T}|\zeta(1/2+{\rm i}t)|^{2k}{\rm d}t$
choosing $A$ large enough and $b_{2}=1/3$. Combining with (74), (77), and (78), we arrive at the claimed statement.
Appendix A A Mollification Lemma
This section contains a mollification lemma that was first developed by one of the authors together with Bourgade and Radziwiłł in conjunction with the work [ABR23]. The lemma was in the end no longer needed in the proof and was therefore unpublished. We include it here as we believe that it is of independent interest. The content of the lemma is to show that $\zeta$ can be well approximated by $\mathcal{M}^{-1}$ for some appropriate mollifier $\mathcal{M}$. The error of the approximation can be made smaller if this approximation is made on an event represented by a Dirichlet polynomial $\mathcal{Q}$. In that regard, the lemma can be seen as a substantial extension of Lemma 4.2 of [ABB+19], with additional technical difficulties coming from the presence of the arbitrary Dirichlet polynomial $\mathcal{Q}$. To state the lemma, we first recall the following definition from [ABR20].
We say a Dirichlet polynomial $\mathcal{Q}$ is $-\mathcal{L}$ well-factorable if it can be written as
$\prod_{0\leq\ell\leq\mathcal{L}}\mathcal{Q}_{\ell}(s)$
where
$\mathcal{Q}_{\ell}(s)=\sum_{\begin{subarray}{c}p|m\Rightarrow p\in(T_{\ell-1},T_{\ell}]\\ \Omega_{\ell}(k)\leq\nu_{\ell}\end{subarray}}\frac{\gamma_{\ell}(k)}{k^{s}},\quad\nu_{\ell}=\frac{1}{10}\mu_{\ell}$
and the coefficients $\gamma_{\ell}(k)$ are arbitrary. Here, $\Omega_{\ell}(k)$ is the number of prime factors of $k$ in the interval $(T_{\ell-1},T_{\ell}]$, and $\mu_{\ell}$ defines the length of the mollifier in (29). Note that at $\ell=0$, there are no restrictions on the number of prime factors in $\mathcal{M}_{0}$ given in (9), so $\mu_{0}=\infty$ effectively.
A $-\ell$ well-factorable Dirichlet polynomial $\mathcal{Q}(s)=\sum_{k}\frac{\gamma(k)}{k^{s}}$ is short for our purpose since the length is
(79) $\leq\prod_{\ell\leq\mathcal{L}}T_{\ell}^{\nu_{\ell}}\leq\exp\Big{(}e^{t}\sum_{\ell\leq\mathcal{L}}\frac{10(\alpha\vee 1)\mathfrak{s}\log_{\ell-1}t)}{(\log_{\ell-1}t)^{\mathfrak{s}}}\Big{)}\leq\exp\Big{(}e^{t}20(\alpha\vee 1)\mathfrak{s}(\log_{\mathcal{L}-1}t)^{1-\mathfrak{s}}\Big{)}\leq T^{1/100},$
for any $\ell\leq\mathcal{L}$ by the choice of $\mathfrak{s}$ in (12).
Lemma 5.
Consider the mollifiers $\mathcal{M}_{\ell}$, $\ell\leq\mathcal{L}$ (see (29) and (31)) and write
$\sigma_{\mathcal{L}}=\frac{1}{2}+\frac{\delta}{e^{t_{\mathcal{L}}}}$
for some fixed $\delta>0$. For any $-\mathcal{L}$ well-factorable $\mathcal{Q}$, we have for $T$ large enough
$\mathbf{E}\left[\Big{|}\zeta\mathcal{M}-1\Big{|}^{2}|\mathcal{Q}|^{2}(\sigma_{\mathcal{L}}+{\rm i}\tau)\right]\leq\Big{(}\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}/10}\Big{)}\cdot\mathbf{E}\Big{[}|\mathcal{Q}|^{2}(\sigma_{\mathcal{L}}+{\rm i}\tau)\Big{]}.$
Proof.
To lighten the notation, we write throughout the proof $\sigma=\sigma_{\mathcal{L}}$ and recall that $\mathcal{M}=\prod_{0\leq\ell\leq\mathcal{L}}\mathcal{M}_{\ell}$. We also sometimes drop the dependence on $\sigma_{\mathcal{L}}+{\rm i}\tau$ in the notation. The proof is by evaluating the expectations for each term in the expansion
(80) $|\zeta\mathcal{M}-1|^{2}=|\zeta\mathcal{M}|^{2}-\zeta\mathcal{M}-\overline{\zeta\mathcal{M}}+1.$
Estimating $\mathbf{E}[\zeta\mathcal{M}|\mathcal{Q}|^{2}]$.
We prove
(81) $\mathbf{E}[\zeta\mathcal{M}|\mathcal{Q}|^{2}]=\mathbf{E}[|\mathcal{Q}|^{2}]\cdot(1+\operatorname{O}(T^{-1/2})).$
The same obviously holds for $\mathbf{E}[\overline{\zeta\mathcal{M}}|\mathcal{Q}|^{2}]$. We use the approximation (see for example Theorem 1.8 in [Ivi85])
(82) $\zeta(\sigma+{\rm i}\tau)=\sum_{n\leq T}\frac{1}{n^{\sigma+{\rm i}\tau}}+\operatorname{O}(T^{-1/2}).$
We define $a_{\ell}(n)=1$ if $\Omega_{\ell}(n)\leq\mu_{\ell}$ and all prime factors of $n$ are in $(T_{\ell-1},T_{\ell}]$, and 00 otherwise. For $n\in\mathbb{N}$, we also write $n_{\ell}=\prod_{p\mid n,p\in(T_{\ell-1},T_{\ell}]}p^{v_{p}(n)}$, where $v_{p}(n)$ stands for the $p$-valuation of $n$. We then get a multiplicative function $a(n)=\prod_{\ell\leq\mathcal{L}}a_{\ell}(n_{\ell})$. As $\mu$ is also multiplicative, we can now write the mollifier in a compact form
$\mathcal{M}(s)=\sum_{m}\frac{\mu(m)a(m)}{m^{s}}.$
By re-defining the coefficient $\gamma$ in the same way to include the restriction on prime factors, we can also write $\mathcal{Q}(s)=\sum_{k}\frac{\gamma(k)}{k^{s}},$ so that
(83) $\mathcal{M}|\mathcal{Q}|^{2}(s)=\sum_{m,k_{1}k_{2}}\frac{\mu(m)a(m)\gamma(k_{1})\overline{\gamma}(k_{2})}{(mk_{1}k_{2})^{s}}.$
The contribution from the error term in (82) is
(84) $\ll T^{-1/2}\sum_{m,k_{1},k_{2}}\frac{|\gamma(k_{1})\gamma(k_{2})|}{(mk_{1}k_{2})^{\sigma}}\ll T^{-\frac{1}{2}+\frac{1}{100}}\left(\sum_{k}\frac{|\gamma(k)|}{k^{\sigma}}\right)^{2}\ll T^{-\frac{1}{2}+\frac{2}{100}}\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}},$
where we have used the bound (32) on $m$ in the second inequality, and the Cauchy-Schwarz inequality in the third. We now observe that $\mathbf{E}[|\mathcal{Q}|^{2}]$ is close to $\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}$ since
$\mathbf{E}[|\mathcal{Q}|^{2}]-\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}=\sum_{k_{1}\neq k_{2}}\frac{\gamma(k_{1})\overline{\gamma}(k_{2})}{(k_{1}k_{2})^{\sigma}}\mathbf{E}\left[\left(\frac{k_{2}}{k_{1}}\right)^{{\rm i}\tau}\right]$ $\ll\sum_{k_{1},k_{2}}\frac{|\gamma(k_{1})\gamma(k_{2})|}{(k_{1}k_{2})^{\sigma}}\frac{k_{2}}{T}$
$\ll T^{-1+\frac{1}{100}}\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}.$
We used $\mathbf{E}[(\frac{k_{2}}{k_{1}})^{{\rm i}\tau}]\ll 1/(T|\log(k_{1}/k_{2})|)\ll\frac{k_{2}}{T}$, (79) to bound $k_{2}$, and the Cauchy-Schwarz inequality. This shows that
(85) $\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}=(1+\operatorname{O}(T^{-1/2}))\mathbf{E}[|\mathcal{Q}|^{2}].$
The above equation together with (84) then yields
(86) $\mathbf{E}\Big{[}\zeta\mathcal{M}|\mathcal{Q}|^{2}(\sigma+{\rm i}\tau)\Big{]}=\sum_{n\leq T,m,k_{1}k_{2}}\frac{\mu(m)a(m)\gamma(k_{1})\overline{\gamma}(k_{2})}{(nmk_{1}k_{2})^{\sigma}}\mathbf{E}\left[\Big{(}\frac{k_{2}}{nmk_{1}}\Big{)}^{{\rm i}\tau}\right]+\operatorname{O}(T^{-1/2})\mathbf{E}[|\mathcal{Q}|^{2}].$
We first bound the terms $nm\neq k_{2}/k_{1}$, based on $\mathbf{E}[(\frac{k_{2}}{nmk_{1}})^{{\rm i}\tau}]\ll 1/(T|\log\frac{nmk_{1}}{k_{2}}|)\ll\frac{k_{2}}{T}$, and (79) and (32) as before:
(87) $\sum_{nm\neq k_{2}/k_{1}}\frac{\mu(m)a(m)\gamma(k_{1})\overline{\gamma}(k_{2})}{(mn)^{\sigma}}\mathbf{E}\left[\left(\frac{k_{2}}{k_{1}nm}\right)^{{\rm i}\tau}\right]$ $\ll T^{\frac{1}{100}}\sum_{n\leq T,k_{1},k_{2}}\frac{\gamma(k_{1})\overline{\gamma}(k_{2})}{(mn)^{\sigma}}\frac{k_{2}}{T}$
$\ll T^{-1+\frac{2}{100}}\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}$
$\ll T^{-1+\frac{2}{100}}\cdot\mathbf{E}[|\mathcal{Q}|^{2}],$
where (85) was used for the last bound.
It remains to estimate the contribution of the terms $nm=k_{2}/k_{1}$. Note that the restriction $n\leq T$ is no longer needed here since it is implied by $n=k_{2}/(k_{1}m)$. Moreover, $m=k_{2}/(k_{1}n)$ and $\Omega_{\ell}(k_{2})\leq\mu_{\ell}$ imposes $a(m)=1$. With these simple observations, we are left to estimate, for fixed $k_{1}$ dividing $k_{2}$, the sum $\sum_{nm=k_{2}/k_{1}}\mu(m)=\mathbf{1}_{k_{2}=k_{1}}.$ With Equations (85), (86) and (87), this gives (81).
Estimating $\mathbf{E}[|\zeta\mathcal{M}\mathcal{Q}|^{2}]$.
We will now prove the more delicate asymptotics
(88) $\mathbf{E}[|\zeta\mathcal{M}\mathcal{Q}|^{2}]=\mathbf{E}[|\mathcal{Q}|^{2}]\cdot\left(1+\operatorname{O}\Big{(}\frac{e^{-2\delta}}{\delta}+e^{-\mu_{\mathcal{L}}}\Big{)}\right),$
relying on the expansion
(89) $\mathbf{E}[|\zeta\mathcal{M}\mathcal{Q}|^{2}]=\sum_{k_{1},k_{2},m_{1},m_{2}}\frac{\mu(m_{1})\mu(m_{2})a(m_{1})a(m_{2})\gamma(k_{1})\overline{\gamma(k_{2})}}{(m_{1}m_{2}k_{1}k_{2})^{\sigma}}\mathbf{E}\left[|\zeta|^{2}\left(\frac{k_{1}m_{1}}{k_{2}m_{2}}\right)^{{\rm i}\tau}\right],$
and on the estimate (see [RS17, Lemma 4])
$\int_{T}^{2T}\Big{(}\frac{n_{1}}{n_{2}}\Big{)}^{{\rm i}t}\left|\zeta\left(\sigma+{\rm i}t\right)\right|^{2}{\rm d}t=\int_{T}^{2T}$ $\Big{(}\zeta(2\sigma)\Big{(}\frac{(n_{1},n_{2})^{2}}{n_{1}n_{2}}\Big{)}^{\sigma}+\Big{(}\frac{t}{2\pi}\Big{)}^{1-2\sigma}\zeta(2-2\sigma)\Big{(}\frac{(n_{1},n_{2})^{2}}{n_{1}n_{2}}\Big{)}^{1-\sigma}\Big{)}{\rm d}t$
(90) $+\operatorname{O}(T^{1-\sigma+\varepsilon}\min(n_{1},n_{2})).$
As $k_{1},m_{1},k_{2},m_{2}\leq T^{1/100}$, and $|a(m_{1})|,|a(m_{2})|\leq 1$, the contribution in (89) from the error term in (90) is
$\ll T^{-\frac{1}{4}}\sum_{k_{1},k_{2}}\frac{|\gamma(k_{1})\gamma(k_{2})|}{(k_{1}k_{2})^{\sigma}}\ll T^{-\frac{1}{10}}\sum_{k}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}\leq T^{-\frac{1}{100}}\cdot\mathbf{E}[|\mathcal{Q}|^{2}],$
where we have used (85) in the last inequality.
Contribution from the right of ${\rm Re}\ s=1$
We now consider the contribution to (89) from the term $\zeta(2\sigma)\Big{(}\frac{(h,k)^{2}}{hk}\Big{)}^{\sigma}$ in (90):
(91) $\zeta(2\sigma)\sum_{k_{1},k_{2}}\frac{\gamma(k_{1})\overline{\gamma(k_{2})}}{(k_{1}k_{2})^{2\sigma}}\sum_{m_{1},m_{2}}\frac{\mu(m_{1})\mu(m_{2})a(m_{1})a(m_{2})}{(m_{1}m_{2})^{2\sigma}}(k_{1}m_{1},k_{2}m_{2})^{2\sigma}.$
We first notice that we can estimate the sums for each interval of primes $(T_{\ell-1},T_{\ell}]$ and take the product on $\ell$ at the very end, by the definition of the $a$’s and the fact that $\mathcal{Q}$ is well-factorable.
With this in mind, let’s first estimate the sum without the restriction on $\Omega_{\ell}$. Fix $k_{1}$ and $k_{2}$. Assume first that $k_{1}\neq k_{2}$. Without loss of generality, we can assume that $(k_{1},k_{2})=1$, otherwise we can factor it out. We write $\mathcal{P}_{1}$ for the set of prime factors of $k_{1}$ and similarly for $k_{2}$. The sum over $m_{1}$ and $m_{2}$ without the restriction can then be expressed as an Euler product as follows
$\sum_{\begin{subarray}{c}m_{1},m_{2}\\ p|m_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2\end{subarray}}\frac{\mu(m_{1})\mu(m_{2})}{(m_{1}m_{2})^{2\sigma}}(k_{1}m_{1},k_{2}m_{2})^{2\sigma}$
(92) $=\prod_{p\in(\mathcal{P}_{1}\cup\mathcal{P}_{2})^{c}}(1-p^{-2\sigma}-p^{-2\sigma}+p^{-2\sigma})\prod_{p\in\mathcal{P}_{1}}(1-1+p^{-2\sigma}-p^{-2\sigma})\prod_{p\in\mathcal{P}_{2}}(1-1+p^{-2\sigma}-p^{-2\sigma}).$
This is obviously 00. With this crucial observation, we have that
$\sum_{p|k_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2}\frac{\gamma(k_{1})\overline{\gamma(k_{2})}}{(k_{1}k_{2})^{2\sigma}}$ $\sum_{p|m_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2}\frac{\mu(m_{1})\mu(m_{2})}{(m_{1}m_{2})^{2\sigma}}(k_{1}m_{1},k_{2}m_{2})^{2\sigma}$
$=\sum_{p|k\Rightarrow p\in(T_{\ell-1},T_{\ell}]}\frac{|\gamma(k)|^{2}}{k^{2\sigma}}\sum_{p|m_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2}\frac{\mu(m_{1})\mu(m_{2})}{(m_{1}m_{2})^{2\sigma}}(m_{1},m_{2})^{2\sigma}$
(93) $=(1+\operatorname{O}(T^{-1/2}))\mathbf{E}[|\mathcal{Q}_{\ell}|^{2}]\prod_{p\in(T_{\ell-1},T_{\ell}]}\left(1-\frac{1}{p^{2\sigma}}\right),$
where we used (85).
We now bound the contribution from integers $m_{1}$ with $\Omega_{\ell}(m_{1})>\mu_{\ell}$ for a fixed $\ell\geq 1$. We use Rankin’s trick, i.e., $\mathbf{1}(\Omega_{\ell}(m)>\mu_{\ell})\leq e^{-\mu_{\ell}}e^{\Omega_{\ell}(m)}$. We get that the sum is bounded by
(94) $e^{-\mu_{\ell}}\sum_{\begin{subarray}{c}m_{1},m_{2},k_{1},k_{2}\\ p|k_{i},m_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2\end{subarray}}\frac{\gamma(k_{1})\overline{\gamma(k_{2})}}{(k_{1}k_{2})^{2\sigma}}\cdot\frac{|\mu(m_{1})\mu(m_{2})|e^{\Omega_{\ell}(m_{1})}}{(m_{1}m_{2})^{2\sigma}}(k_{1}m_{1},k_{2}m_{2})^{2\sigma}.$
The sum over $m_{1},m_{2}$ is computed similarly to (92). It is
$=(k_{1},k_{2})^{2\sigma}\prod_{\mathcal{P}_{1}}\left(2+\frac{2e}{p^{2\sigma}}\right)\prod_{\mathcal{P}_{2}}\left(1+e+\frac{1+e}{p^{2\sigma}}\right)\prod_{(\mathcal{P}_{1}\cup\mathcal{P}_{2})^{c}}\left(1+\frac{2e+1}{p^{2\sigma}}\right)$
$\leq(k_{1},k_{2})^{2\sigma}(1+e)^{\omega_{\ell}(k_{1})+\omega_{\ell}(k_{2})}\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)},$
where $\omega_{\ell}(m)$ stands for the number of distinct prime factors of $m$ in $(T_{\ell-1},T_{\ell}]$. Using the fact that $|\gamma(k_{1})\gamma(k_{2})|\leq(|\gamma(k_{1})|^{2}+|\gamma(k_{2})|^{2})/2$, the sum in (94) is bounded by
(95) $\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)}\sum_{\begin{subarray}{c}k_{1}\\ p|k_{1}\Rightarrow p\in(T_{\ell-1},T_{\ell}]\end{subarray}}\frac{|\gamma(k_{1})|^{2}}{k_{1}^{2\sigma}}(1+e)^{\omega_{\ell}(k_{1})}\sum_{\begin{subarray}{c}k_{2}\\ p|k_{2}\Rightarrow p\in(T_{\ell-1},T_{\ell}]\end{subarray}}\frac{(k_{1},k_{2})^{2\sigma}}{k_{2}^{2\sigma}}(1+e)^{\omega_{\ell}(k_{2})}.$
For $k_{1}$ fixed, we can express the sum over $k_{2}$ as an Euler product
$=\prod_{p|k_{1}}\Big{(}1+(1+e)v_{p}(k_{1})+(1+e)\sum_{j>v_{p}(k_{1})}p^{-2\sigma(j-v_{p}(k_{1}))}\Big{)}\prod_{p\nmid k_{1}}\Big{(}1+(1+e)\sum_{j\geq 1}p^{-2\sigma j}\Big{)}$
$\leq(1+e)^{\omega_{\ell}(k_{1})}\prod_{p\mid k_{1}}(1+v_{p}(k_{1}))\prod_{p\nmid k_{1}}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)}$
$\leq(1+e)^{\Omega_{\ell}(k_{1})}e^{\Omega_{\ell}(k_{1})}\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)},$
where we use the fact that $1+x\leq e^{x}$ to estimate the product over $p|k_{1}$. Using the assumption that $\Omega_{\ell}(k_{1})\leq\nu_{\ell}$, we can finally bound (94) by
(96) $\leq(1+\operatorname{O}(T^{-1/2}))(1+e)^{3\nu_{\ell}-\mu_{\ell}}\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)}^{2}\mathbf{E}[|\mathcal{Q}_{\ell}|^{2}],$
where we used (85). By the choice of $\nu_{\ell}$, we have that $23\nu_{\ell}-\mu_{\ell}<\mu_{\ell}/$ 2.
To estimate (91), it remains to take the product over $\ell$ and multiply by ${\zeta(2\sigma)=\prod_{p}(1-p^{-2\sigma})^{-1}}$. We now apply the identity
$\prod_{\ell\leq\mathcal{L}}(1-\mathbf{1}(\Omega_{\ell}(m)>\mu_{\ell}))=\sum_{I\subseteq\{1,\dots,\mathcal{L}\}}\prod_{\ell\in I}(-1)^{|I|}\mathbf{1}(\Omega_{\ell}(m)>\mu_{\ell})$
to the sums in $m_{1}$ and $m_{2}$ in (91) to get that it is equal to
(97) $\mathbf{E}[|\mathcal{Q}|^{2}]\Big{\{}(1+\operatorname{O}(T^{-1/2}))\prod_{p>T_{\mathcal{L}}}\left(1-\frac{1}{p^{2\sigma}}\right)^{-1}-\\ \operatorname{O}^{\star}\Big{(}\sum_{I_{1}\cup I_{2}\neq\emptyset}\prod_{\ell\in I_{1}\cup I_{2}}e^{-\mu_{\ell}/2}\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)}^{4}\Big{(}1-\frac{1}{p^{2\sigma}}\Big{)}^{-1}\Big{)}\Big{\}}.$
Here the sums is over subsets $I_{1}$ and $I_{2}$ whose union is not empty. They represent the $\ell$’s for which $\Omega_{\ell}(m)>\mu_{\ell}$ for $m_{1}$ and $m_{2}$ respectively. As in Lemma 6, $\operatorname{O}^{\star}$ represents an error term with implicit constant less than $1$. We then use the estimate (93) when $\ell\in(I_{1}\cup I_{2})^{c}$ and the estimate (96) when $\ell\in I_{1}\cup I_{2}$. Note that in the case where $\Omega_{\ell}(m)>\mu_{\ell}$ for both $m_{1}$ and $m_{2}$, we need to square the term $\Big{(}1+\frac{2e+1}{p^{2\sigma}}\Big{)}^{2}$. The first product in the braces is estimated by the Prime Number Theorem:
$\sum_{p>T_{\mathcal{L}}}\frac{1}{p^{2\sigma}}\leq\int_{T_{\mathcal{L}}}^{\infty}\frac{1}{u^{2\sigma}\log u}{\rm d}u+\operatorname{O}(e^{-c\sqrt{T_{\mathcal{L}}}})$ $=\int_{\log T_{\mathcal{L}}}^{\infty}\frac{e^{-v(2\sigma-1)}}{v}{\rm d}v+\operatorname{O}(e^{-c\sqrt{T_{\mathcal{L}}}})$
$\leq\frac{e^{-2\delta}}{2\delta}+\operatorname{O}(e^{-c\sqrt{T_{\mathcal{L}}}}),$
by the definition of $\sigma$ and $\delta$. Therefore, $\prod_{p>T_{\mathcal{L}}}\left(1-p^{-2\sigma}\right)^{-1}$ is $1+e^{-2\delta}/\delta$ for $T$ large enough.
Write $\rho_{\ell}=\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1+(2e+1)p^{-2\sigma}\Big{)}^{4}\Big{(}1-p^{-2\sigma}\Big{)}^{-1}$. We have that locally the factors are bounded by $(1+100p^{-2\sigma})$. Moreover, Mertens’s theorem implies $e^{-\mu_{\ell}/2}\rho_{\ell}\leq e^{-\mu_{\ell}/4}$ (for $\mathfrak{s}>100$). The $\operatorname{O}^{\star}$-term is then
$\sum_{I_{1}\cup I_{2}\neq\emptyset}\prod_{\ell\in I_{1}\cup I_{2}}e^{-\mu_{\ell}/2}\rho_{\ell}\leq\prod_{\ell\leq\mathcal{L}}(1+e^{-\mu_{\ell}/4})^{2}-1\leq e^{-\mu_{\ell}/10}.$
This concludes the proof of the contribution in (89) from the term $\zeta(2\sigma)\Big{(}\frac{(n_{1},n_{2})^{2}}{n_{1}n_{2}}\Big{)}^{\sigma}$ in (90).
Contribution from the left of ${\rm Re}\ s=1$
We now consider the contribution from $\Big{(}\frac{t}{2\pi}\Big{)}^{1-2\sigma}\zeta(2-2\sigma)\Big{(}\frac{(n_{1},n_{2})^{2}}{n_{1}n_{2}}\Big{)}^{1-\sigma}$ in (90). The integral factor there is
(98) $\frac{1}{T}\int_{T}^{2T}\left(\frac{t}{2\pi}\right)^{1-2\sigma}{\rm d}t\ll T^{1-2\sigma}=\exp(-2\delta e^{t-t_{\mathcal{L}}}).$
For the other terms, we can work on each interval $(T_{\ell-1},T_{\ell}]$ for $\ell$ fixed as before. Without the restriction on the number of prime factors, the sum over $k$’s and $m$’s is
(99) $\sum_{\begin{subarray}{c}m_{1},m_{2},k_{1},k_{2}\\ p|k_{i},m_{i}\Rightarrow p\in(T_{\ell-1},T_{\ell}],i=1,2\end{subarray}}\frac{\gamma(k_{1})\overline{\gamma(k_{2})}}{k_{1}k_{2}}\frac{\mu(m_{1})\mu(m_{2})}{m_{1}m_{2}}(k_{1}m_{1},k_{2}m_{2})^{2-2\sigma}.$
We write $k_{1}=(k_{1},k_{2})l_{1}$, $k_{2}=(k_{1},k_{2})l_{2}$ where $l_{1}$ and $l_{2}$ are coprime with corresponding prime divisors sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. The above sum over $m_{1},m_{2}$ is equal to
$\sum_{\begin{subarray}{c}m_{1},m_{2}\\ p|m_{1},m_{2}\Rightarrow p\in(T_{\ell-1},T_{\ell}]\end{subarray}}\frac{\mu(m_{1})\mu(m_{2})}{m_{1}m_{2}}(k_{1}m_{1},k_{2}m_{2})^{2-2\sigma}$
(100) $\quad=(k_{1},k_{2})^{2-2\sigma}\prod_{p\in(\mathcal{P}_{1}\cup\mathcal{P}_{2})^{c}}(1-2p^{-1}+p^{-2\sigma})\prod_{p\in\mathcal{P}_{1}}(1-p^{-1}-p^{1-2\sigma}+p^{-2\sigma})\prod_{p\in\mathcal{P}_{2}}(1-p^{-1}-p^{1-2\sigma}+p^{-2\sigma}).$
As opposed to (92), the above product does not vanish when $k_{1}\neq k_{2}$. Therefore, estimating the sum is slightly more involved. We define
$c_{\ell}=(2\sigma-1)\log T_{\ell}.$
On $(\mathcal{P}_{1}\cup\mathcal{P}_{2})^{\rm c}$, we bound
$1-p^{-1}-p^{-1}+p^{-2\sigma}\leq 1-p^{-1},$
and on $\mathcal{P}_{1}\cup\mathcal{P}_{2}$, we have when $p\leq T_{\ell}$,
$1-p^{-1}-p^{1-2\sigma}+p^{-2\sigma}=(1-p^{-1})\cdot(1-p^{1-2\sigma})\leq(1-p^{-1})c_{\ell}.$
This shows that the sum in (100) is
(101) $\leq(k_{1},k_{2})^{2-2\sigma}\prod_{p\in(T_{\ell-1},T_{\ell}]}\left(1-\frac{1}{p}\right)\cdot c_{\ell}^{\omega_{\ell}(k_{1})+\omega_{\ell}(k_{2})}.$
Moreover, bounding the $\gamma$’s as before, we get that (100) is
$\leq\sum_{\begin{subarray}{c}k_{1},k_{2}\\ p|k_{1},k_{2}\Rightarrow p\in(T_{\ell-1},T_{\ell}]\end{subarray}}\frac{|\gamma(k_{1})|^{2}}{k_{1}k_{2}}(k_{1},k_{2})^{2-2\sigma}c_{\ell}^{\omega_{\ell}(k_{1})+\omega_{\ell}(k_{2})}.$
The sum over $k_{2}$ can be evaluated for fixed $k_{1}$ by re-expressing it as an Euler product. We get that
$\sum_{\begin{subarray}{c}k_{2}\\ p|k_{2}\Rightarrow p\in(T_{\ell-1},T_{\ell}]\end{subarray}}$ $c_{\ell}^{\omega_{\ell}(k_{2})}\frac{(k_{1},k_{2})^{2-2\sigma}}{k_{2}}$
$=\prod_{p\nmid k_{1}}\Big{(}1+\sum_{j=1}^{v_{p}(k_{1})}c_{\ell}p^{j(1-2\sigma)}+\sum_{j>v_{p}(k_{1})}c_{\ell}p^{v_{p}(k_{1})(2-2\sigma)}{p^{-j}}\Big{)}\prod_{p\mid k_{1}}\Big{(}1+\sum_{j\geq 1}\frac{c_{\ell}}{p^{j}}\Big{)}$
$\leq\prod_{p\mid k_{1}}\Big{(}1+c_{\ell}v_{p}(k_{1})p^{1-2\sigma}+c_{\ell}p^{v_{p}(k_{1})(1-2\sigma)}\Big{)}\prod_{p\nmid k_{1}}\Big{(}1+\frac{c_{\ell}}{p-1}\Big{)}.$
We factor $p^{(1-2\sigma)v_{p}(k_{1})}$ from the first product to get $k_{1}^{1-2\sigma}$. Combined with $k_{1}$, this will get the factor $k_{1}^{-2\sigma}$ needed to recover $\mathcal{Q}_{\ell}^{2}$. The second product is bounded by $\exp(2c_{\ell}(t_{\ell}-t_{\ell-1}))$ by Mertens’s theorem. Putting this together yields the bound
$k_{1}^{1-2\sigma}e^{2c_{\ell}(t_{\ell}-t_{\ell-1})}\prod_{p\mid k_{1}}\Big{(}p^{v_{p}(k_{1})(2\sigma-1)}+c_{\ell}v_{p}(k_{1})p^{(v_{p}(k_{1})-1)(2\sigma-1)}+c_{\ell}\Big{)}$
$\leq k_{1}^{1-2\sigma}e^{2c_{\ell}(t_{\ell}-t_{\ell-1})}(c_{\ell}+c_{\ell}\nu_{\ell}e^{c_{\ell}\nu_{\ell}}+e^{c_{\ell}\nu_{\ell}})^{\nu_{\ell}}\leq k_{1}^{1-2\sigma}e^{2c_{\ell}(t_{\ell}-t_{\ell-1})+c_{\ell}\nu_{\ell}^{2}}.$
where we used the definition of $c_{\ell}$ and the bound $\Omega_{\ell}(k_{1})\leq\mu_{\ell}$ to get the inequality. Taking the product over $\ell$ and including the integral factor (98), we can now conclude that the contribution to (99) of sums on $m$’s and $k$’s without the restriction on the number of factors of the $m$’s is
$\zeta(2-2\sigma)\prod_{p\leq T_{\mathcal{L}}}(1-p^{-1})e^{\sum_{\ell\leq\mathcal{L}}2c_{\ell}\nu_{\ell}^{2}-2\delta e^{t-t_{\mathcal{L}}}}\sum_{k_{1}}\frac{|\gamma(k_{1})|^{2}}{k_{1}^{2\sigma}}.$
The first two factors are of order one. The sum over $k_{1}$ is $\ll\mathbf{E}[|\mathcal{Q}|^{2}]$ by (85). Since $c_{\mathcal{L}}=2\delta$, the exponential factor is
$\leq\exp\Big{(}4c_{\mathcal{L}}\nu_{\mathcal{L}}^{2}-2\delta e^{t-t_{\mathcal{L}}}\Big{)}\leq\exp\Big{(}8\delta(10\mathfrak{s}\log_{\mathcal{L}}t)^{2}-2\delta e^{\mathfrak{s}\log_{\mathcal{L}}t}\Big{)}\leq e^{-4\delta},$
under the condition that $\mathfrak{s}\log_{\mathcal{L}}t>100$. This is the dominant contribution coming from ${\rm Re}s=2-2\sigma$. Finally, we need to bound the contribution to the sum over $m$’s of the restrictions $\Omega_{\ell}(m)>\mu_{\ell}$ for each $\ell$. This is done exactly as before starting from (94). ∎
Appendix B Auxiliary Results
The following lemma is an adaption of Lemma 23 in [ABR23] that gives upper and lower bound in approximating $\mathcal{M}_{\ell}$ by $e^{-(S_{\ell}-S_{\ell-1})}$ with precise error.
Lemma 6.
Consider the partial sums (73). Let $\ell\geq 1$ and $A>0$. If $|\widetilde{S}_{\ell}-\widetilde{S}_{\ell-1}|\leq A(t_{\ell}-t_{\ell-1})$, then the mollifier $\mathcal{M}_{\ell}$ defined in (29) is satisfies
$|\mathcal{M}_{\ell}|=\Big{(}1+\operatorname{O}^{\star}\big{(}\tfrac{5}{\sqrt{T_{\ell-1}}}\big{)}\Big{)}e^{-(S_{\ell}-S_{\ell-1})}+e^{-\mu_{\ell}+5(A+1)(t_{\ell}-t_{\ell-1})+1},$
where $\operatorname{O}^{\star}$ stands for $\operatorname{O}$ with implicit constant smaller than $1$.
Proof.
We set $s=\sigma+{\rm i}\tau$. We have by expanding the logarithm
$\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1-\frac{1}{p^{s}}\Big{)}$ $=e^{-(S_{\ell}-S_{{\ell-1}})-R_{\ell}},$
where
$R_{\ell}=\sum_{a\geq 3}\sum_{p\in(T_{\ell-1},T_{\ell}]}\frac{1}{a}\,{\rm Re}\leavevmode\nobreak\ p^{-as}.$
We estimate $R_{\ell}$ as follows
$|R_{\ell}|$ $\leq\sum_{a\geq 3}\sum_{p\in(T_{\ell-1},T_{\ell}]}\frac{1}{a}p^{-a/2}\leq\frac{1}{3}\frac{\sqrt{2}}{\sqrt{2}-1}\int_{T_{\ell-1}}^{\infty}y^{-3/2}{\rm d}y\leq 2.5(T_{\ell-1})^{-1/2}.$
Using the fact that $e^{x}\leq 1+2x$ for $20<x<1/$ 2 and $e^{-x}\geq 1-x$ for $x<0$, we get that
(102) $1-(T_{\ell-1})^{-1/2}\leq e^{-R_{\ell}}\leq 1+5(T_{\ell-1})^{-1/2}.$
It remains to relate the product over primes to the mollifier. By definition we have
$\prod_{p\in(T_{\ell-1},T_{\ell}]}\Big{(}1-\frac{1}{p^{s}}\Big{)}$ $=\mathcal{M}_{\ell}+\sum_{\begin{subarray}{c}p|n\Rightarrow p\in(T_{\ell-1},T_{\ell}]\\ \Omega_{\ell-1}(n)>\mu_{\ell}\end{subarray}}\frac{\mu(n)}{n^{s}}.$
The last term is equal to
(103) $\sum_{j>\mu_{\ell}}(-1)^{j}\Big{(}\sum_{T_{\ell-1}<p_{1}<\ldots<p_{j}\leq T_{\ell}}\frac{1}{(p_{1}\ldots p_{j})^{s}}\Big{)}.$
The Girard-Newton identities (see for example Equation 2.14’ in [Mac95]) show that the inner sum is
$\sum_{T_{\ell-1}<p_{1}<\ldots<p_{j}\leq T_{\ell}}\frac{1}{(p_{1}\ldots p_{j})^{s}}=(-1)^{j}\sum_{\begin{subarray}{c}m_{1},\ldots,m_{k},\ldots\geq 0\\ m_{1}+2m_{2}+\ldots+km_{k}+(k+1)m_{k+1}+\ldots=j\end{subarray}}\prod_{k\geq 1}\frac{(-\mathcal{P}(ks))^{m_{k}}}{m_{k}!\ k^{m_{k}}}$
where the sum is over the partitions of $j$ and
$\mathcal{P}(s)=\sum_{T_{\ell-1}<p\leq T_{\ell}}\frac{1}{p^{s}}.$
Using this we can bound the absolute value of (103) by
$\sum_{m_{1},\ldots,m_{k},\ldots\geq 0}$ $\exp\Big{(}-\mu_{\ell}+m_{1}+2m_{2}+\ldots+km_{k}+\ldots\Big{)}\prod_{k\geq 1}\frac{|\mathcal{P}(ks)|^{m_{k}}}{m_{k}!\ k^{m_{k}}}$
(104) $=e^{-\mu_{\ell}}\prod_{k\geq 1}\Big{(}\sum_{m_{k}\geq 0}\frac{(e^{k}|\mathcal{P}(ks)|/k)^{m_{k}}}{m_{k}!}\Big{)}$
$=\exp\Big{(}-\mu_{\ell}+\sum_{k\geq 1}e^{k}|\mathcal{P}(ks)|/k\Big{)}.$
For $\ell=1$, the assumption implies $|\mathcal{P}(s)|\leq A(t_{\ell}-t_{\ell-1})$, giving a contribution to the sum of
(105) $\leq eA(t_{\ell}-t_{\ell-1}).$
For $k=2$, Mertens’s theorem (see [Van17] for a precise bound) also implies
$|\mathcal{P}(2s)|\leq t_{\ell}-t_{\ell-1}+\operatorname{O}^{\star}\Big{(}\tfrac{4}{\log^{3}T_{\ell-1}}\Big{)}\leq t_{\ell}-t_{\ell-1}+0.001,$
by the definition of $T_{0}$. This gives an error of
(106) $\tfrac{e^{2}}{2}|\mathcal{P}(2s)|\leq 5(t_{\ell}-t_{\ell-1})+0.01.$
For $k\geq 3$, we have
$|\mathcal{P}(ks)|\leq\sum_{p\in(T_{\ell-1},T_{\ell}]}p^{-k/2}\leq 2(T_{\ell-1})^{-k/2+1}.$
This implies that
(107) $\sum_{k\geq 3}e^{k}|\mathcal{P}(ks)|/k\leq 2T_{0}\sum_{k\geq 3}\Big{(}\frac{e}{T_{0}^{1/2}}\Big{)}^{k}\leq 0.001.$
The claim follows from (102), (105), (106), and (107). ∎
The mean-value theorem for Dirichlet polynomials, see [MV74, Corollary 3], takes the following form when expressed in terms of the random model 51. Let $(\theta_{p},p\text{ prime})$ a sequence of IID random variables, uniformly distributed on $[0,2\pi]$. For an integer $n$ with prime factorization $n=p_{1}^{a_{1}}\ldots p_{k}^{a_{k}}$ with $p_{1},\ldots,p_{k}$ all distinct, consider
(108) $\theta_{n}=\sum_{j=1}^{k}\theta_{p_{j}}^{a_{j}}.$
The orthogonality relation $\mathbf{E}[e^{{\rm i}\theta_{n}}\overline{e^{{\rm i}\theta_{m}}}]=\mathbf{1}_{n=m}$ is straightforward from the definition. This implies that for an arbitrary sequence $a(n)$ of complex numbers,
$\mathbf{E}\Big{[}\Big{|}\sum_{n\leq N}a(n)e^{{\rm i}\theta_{n}}\Big{|}^{2}\Big{]}=\sum_{n\leq N}|a(n)|^{2}.$
The mean-value theorem for Dirichlet polynomials is as follows:
Lemma 7 (Corollary 3 in [MV74]).
Let $(a(n),n\geq 1)$ be any sequence of complex numbers. Then for $N\geq 1$ and $T\geq 1$, we have with the notation above.
$\mathbf{E}\Big{[}\Big{|}\sum_{n\leq N}a(n)n^{{\rm i}\tau}\Big{|}^{2}\Big{]}=\Big{(}1+\operatorname{O}\Big{(}\frac{N}{T}\Big{)}\Big{)}\mathbf{E}\Big{[}\Big{|}\sum_{n\leq N}a(n)e^{{\rm i}\theta_{n}}\Big{|}^{2}\Big{]}.$
We need several lemmas controlling the distribution of the random model (51). The first one compares it explicitly to the Gaussian random variables $\mathcal{Z}_{\ell}$.
Lemma 8.
Consider the increments $\mathcal{Y}_{\ell}=\mathcal{S}_{\ell}-S_{\ell-1}$, $1\leq\ell\leq\mathcal{L}$, defined in (52), and the corresponding Gaussian variables $\mathcal{N}_{\ell}=\mathcal{Z}_{\ell}-\mathcal{Z}_{\ell-1}$ as in (14). There exists a constant $c>0$ such that, for any intervals $A$
$\mathbf{P}\Big{(}\mathcal{Y}_{\ell}\in A\Big{)}=\mathbf{P}\Big{(}\mathcal{N}_{\ell}\in A\Big{)}+\operatorname{O}(e^{-ce^{t_{\ell-1}/2}}).$
Proof.
This follows similarly to [ABR20, Lemma 20], based on the Berry-Esseen estimate as stated in [ABR20, Lemma 19]. The proof is actually more immediate because the covariances of $(\mathcal{Y},\mathcal{Y}^{\prime})$ and $(\mathcal{N},\mathcal{N}^{\prime})$ exactly coincide. ∎
The error $\operatorname{O}(e^{-ce^{t_{\ell-1}/2}})$ is too large for $\ell=1$ for our purpose. In this case, we resort to the weaker estimate given in [ABR20, Lemma 18]:
Lemma 9.
Consider the random variables $\mathcal{S}_{\ell}$ defined in (51). Let $\alpha\geq 1$. Then, for all $\Delta\geq\alpha$ and $|u|\leq 100\alpha\ell$, we have
$\mathbb{P}(\mathcal{S}_{\ell}\in[u,u+\Delta^{-1}])\asymp\frac{1}{\Delta}\cdot\frac{e^{-u^{2}/{2v_{\ell}^{2}}}}{\sqrt{v_{\ell}}}.$
When an exact Gaussian comparison is not needed, it is sometimes useful to bound the moments of the random model by Gaussian ones. The following statement is a direct consequence of the fact that
(109) $\mathbf{E}\left[\exp\Big{(}\lambda\sum_{x<p\leq y}\frac{{\rm Re}\ a(p)X(p)}{p^{\sigma}}\Big{)}\right]=\prod_{x<p\leq y}I_{0}\Big{(}\frac{|a(p)|\lambda}{p^{\sigma}}\Big{)},$
where
(110) $I_{0}(w)=\sum_{k=0}^{\infty}\frac{w^{2k}}{2^{2k}(k!)^{2}}.$
Lemma 10.
Let $\sigma\geq 1/2$ and $1\leq x<y$. For any complex numbers $a(p)$, we have for $k\in\mathbb{N}$,
(111) $\mathbf{E}\left[\Big{(}\sum_{x<p\leq y}\frac{{\rm Re}\ a(p)e^{{\rm i}\theta_{p}}}{p^{\sigma}}\Big{)}^{2k}\right]\leq\frac{(2k!)}{2^{k}k!}\mathfrak{s}(\sigma)^{2k},$
where
(112) $\mathfrak{s}^{2}(\sigma)=\frac{1}{2}\sum_{x<p\leq y}\frac{|a(p)|^{2}}{p^{2\sigma}}.$
In particular, we have for any $\lambda\in\mathbb{R}$,
(113) $\mathbf{E}\left[\exp\Big{(}\lambda\sum_{x<p\leq y}\frac{{\rm Re}\ a(p)e^{i\theta_{p}}}{p^{\sigma}}\Big{)}\right]\leq\exp(\lambda^{2}\mathfrak{s}^{2}(\sigma)/2).$
To bound the moments of actual Dirichlet polynomials, we have the useful Lemma 3 of [Sou09]:
Lemma 11.
Let $2\leq x\leq T$. For any complex numbers $a(p)$ and $k\in\mathbb{N}$ with $x^{k}\leq T/\log T$, we have for $T$ large enough
$\mathbf{E}\left[\Big{|}\sum_{p\leq x}\frac{a(p)}{p^{\sigma+{\rm i}\tau}}\Big{|}^{2k}\right]\ll k!\Big{(}\sum_{p\leq x}\frac{|a(p)|^{2}}{p^{2\sigma}}\Big{)}^{k}.$
Using Stirling, we get by picking $k=\lfloor V^{2}/v_{\ell}^{2}\rfloor$ (assuming that $V$ is not too large so the moment condition is satisfied) that
$\mathbf{P}\Big{(}S_{\ell}-S_{\ell-1}>V\Big{)}\ll\sqrt{\frac{V^{2}}{v_{\ell}^{2}-v_{\ell-1}^{2}}}\exp\Big{(}-\frac{V^{2}}{2(v_{\ell}^{2}-v_{\ell-1}^{2})}\Big{)}.$
The asymptotics of the moments of the random model with all the prime powers present, as in the definition of $\mathcal{Z}_{0}$ in (15), can be calculated explicitly.
Lemma 12.
Let $\delta>0$ and $\sigma=\frac{1}{2}+\frac{\delta}{\log T}$. Let $T_{0}$ such that $\log T_{0}={\rm o}(\log T)$. Then, as $T\to\infty$, we have for any $k>0$
(114) $(\log T_{0})^{-k^{2}}\mathbf{E}\left[\left|\prod_{p\leq T_{0}}\left(1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\right)^{-1}\right|^{2k}\right]\sim e^{\gamma k^{2}}\prod_{p}(1-p^{-1})^{k^{2}}\sum_{m\geq 0}\frac{|d_{k}(p^{m})|^{2}}{p^{2m}},$
where $d_{k}(n)$ is the $k$-divisor function defined through the Taylor series
(115) $(1-x)^{-k}=\sum_{n\geq 0}d_{k}(n)x^{n}=\sum_{n\geq 0}\frac{\Gamma(n+k)}{n!\Gamma(k)}x^{n}.$
(For an integer $k$, $d_{k}(n)$ is the number of ways of writing $n$ as the product of $k$ factors.) The Euler product on the right-hand side of (114) is $a_{k}$, cf. (4).
Proof.
Let $2\leq X<T_{0}$. We will take the limit $T\to\infty$ then $X\to\infty$. We have by expanding the logarithm and the exponential
$\mathbf{E}\left[\left|\prod_{X<p\leq T_{0}}\left(1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\right)^{-1}\right|^{2k}\right]$ $=\prod_{X<p\leq T_{0}}\mathbf{E}\left[\exp\Big{(}-2k{\rm Re}\log(1-e^{{\rm i}\theta_{p}}p^{-\sigma})\Big{)}\right]$
$=\prod_{X<p\leq T_{0}}\mathbf{E}\left[\exp\Big{(}2k\Big{(}\frac{\cos\theta_{p}}{p^{\sigma}}+\frac{\cos 2\theta_{p}}{2p^{2\sigma}}+\operatorname{O}(p^{-3\sigma})\Big{)}\Big{)}\right]$
$=\prod_{X<p\leq T_{0}}\left(1+\frac{k^{2}}{p^{2\sigma}}+\operatorname{O}(p^{-3\sigma})\right)$
$=\exp\left(\sum_{X<p\leq T_{0}}\frac{k^{2}}{p^{2\sigma}}+\operatorname{O}(X^{-1/2})\right).$
We can evaluate the sum over $p$ as in (17) using the Prime Number Theorem. This gives as $T\to\infty$ and $X\to\infty$
$\mathbf{E}\left[\left|\prod_{X<p\leq T_{0}}\left(1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\right)^{-1}\right|^{2k}\right]\sim(\log T_{0})^{k^{2}}(\log X)^{-k^{2}},$
since $\delta\frac{\log T_{0}}{\log T}\to 0$ by assumption. It remains to evaluate $(\log X)^{-k^{2}}\mathbf{E}[|\prod_{p\leq X}(1-e^{{\rm i}\theta_{p}}p^{-\sigma})^{-1}|^{2k}]$ in the limit $T\to\infty$ and $X\to\infty$. Taking $T\to\infty$ simplifies $\sigma\to 1/2$. Hence, writing the random Euler product as a random Dirichlet series yields
$\mathbf{E}\left[\left|\prod_{p\leq X}\left(1-\frac{e^{{\rm i}\theta_{p}}}{p^{\sigma}}\right)^{-k}\right|^{2}\right]\sim\mathbf{E}\Big{[}\Big{|}\sum_{\begin{subarray}{c}n\geq 1\\ p|n\Rightarrow p\leq X\end{subarray}}\frac{d_{k}(n)}{n^{1/2}}e^{{\rm i}\theta_{n}}\Big{|}^{2}\Big{]}$
where $\theta_{n}$ is defined by $\theta_{n}=\theta_{p_{1}}^{a_{1}}\dots\theta_{p_{r}}^{a_{r}}$ if the prime factorization of $n$ is $p_{1}^{a_{1}}\dots p_{r}^{a_{r}}$. The right-hand side is easily evaluated by the orthogonality of the $e^{{\rm i}\theta_{n}}$ and equals
$\sum_{\begin{subarray}{c}n\geq 1\\ p|n\Rightarrow p\leq X\end{subarray}}\frac{|d_{k}(n)|^{2}}{n}=\prod_{p\leq X}\sum_{m\geq 0}\frac{|d_{k}(p^{m})|^{2}}{p^{m}}.$
Mertens’s third theorem, see for example [Kou19], states that $\log X\prod_{p\leq X}(1-p^{-1})\sim e^{-\gamma}$ This implies that as $X\to\infty$
$(\log X)^{-k^{2}}\mathbf{E}\left[\left|\prod_{p\leq X}\left(1-\frac{e^{{\rm i}\theta_{p}}}{p^{1/2}}\right)^{-1}\right|^{2k}\right]\sim e^{\gamma k^{2}}\prod_{p}(1-p^{-1})^{k^{2}}\sum_{m\geq 0}\frac{|d_{k}(p^{m})|^{2}}{p^{m}},$
as claimed. ∎
The lemma implies some useful estimates. First, Markov’s inequality yields
(116) $\mathbf{P}(\mathcal{Z}_{0}>\alpha t_{0})\leq\mathbf{E}[e^{2\alpha\mathcal{Z}_{0}}](\log T_{0})^{-2\alpha^{2}}\ll e^{\gamma\alpha^{2}}a_{\alpha}(\log T_{0})^{-\alpha^{2}}.$
Second, the heuristic given (18) remains almost unchanged on $\mathcal{G}_{0}=\{\mathcal{Z}_{0}-\alpha t_{0}\in[0,\sqrt{t_{0}}\}$. Indeed, let’s consider the change of probability defined by $\frac{{\rm d}\widetilde{\mathbf{P}}}{{\rm d}\mathbf{P}}=\frac{e^{2\alpha\mathcal{Z}_{0}}}{\mathbf{E}[e^{2\alpha\mathcal{Z}_{0}}]}$. We have similar to (18):
$\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t\}\cap\mathcal{G}_{0})$ $\asymp\frac{e^{-\alpha^{2}t}}{\alpha\sqrt{t}}e^{-\alpha^{2}\mathfrak{s}\log_{\mathcal{L}}t-2\delta\alpha^{2}}\cdot e^{-\alpha^{2}t_{0}}\mathbf{E}[e^{2\alpha\mathcal{Z}_{0}}]\cdot\widetilde{\mathbf{P}}(\mathcal{Z}_{0}-\alpha t_{0}\in[0,\sqrt{t_{0}}])$
$\asymp\frac{e^{-\alpha^{2}t}}{\alpha\sqrt{t}}a_{\alpha}e^{-\alpha^{2}\mathfrak{s}\log_{\mathcal{L}}t-2\delta\alpha^{2}+\gamma\alpha^{2}}\cdot\widetilde{\mathbf{P}}(\mathcal{Z}_{0}-\alpha t_{0}\in[0,\sqrt{t_{0}}]),$
by Lemma 12. Under $\widetilde{\mathbf{P}}$, the random variables $e^{{\rm i}\theta_{p}}$ remain independent and $\widetilde{\mathbf{E}}[\mathcal{Z}_{0}]=\alpha t_{0}+{\rm o}(1)$. Therefore, by the Berry-Esseen theorem, we have that $\widetilde{\mathbf{P}}(\mathcal{Z}_{0}-\alpha t_{0}\in[0,\sqrt{t_{0}}])\asymp 1$. We conclude that
(117) $\mathbf{P}(\{\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t\}\cap\mathcal{G}_{0})\asymp\mathbf{P}(\mathcal{Z}_{0}+\mathcal{Z}_{\mathcal{L}}>\alpha t).$